### v3

# Machines are more productive than humans until they aren’t, and vice versa

Abstract
With the growth of artificial skills, organizations are increasingly confronting with the problem of optimizing skill policy decisions guided by economic principles. This paper addresses the underlying complexity of this challenge by developing an in-silico framework based on Monte Carlo simulations grounded in empirical realism to analyze the economic impact of human and machine skills, individually or jointly deployed, in the execution of tasks presenting varying levels of complexity. Our results provide quantitative support for the established notions that automation tends to be the most economically-effective strategy for tasks characterized by low-to-medium generalization difficulty, while automation may struggle to match the economic utility of human skills in more complex scenarios. Critically, our simulations highlight that, when high level of generalization is required and the cost of errors is high, combining human and machine skills can be the most effective strategy, but only if genuine augmentation is achieved. In contrast, when failing to realize this synergy, the human-machine policy is severely penalized by the inherent costs of its dual skill structure, causing it to destroy value and becoming the worst choice from an economic perspective. The takeaway for decision-makers is unambiguous: in complex and critical contexts, simply allocating human and machine skills to a task may be insufficient, and a human-machine skill policy is neither a silver-bullet solution nor a low-risk compromise. Rather, it is a critical opportunity to boost competitiveness that demands a strong organizational commitment to enabling augmentation. Also, our findings show that improving the cost-effectiveness of machine skills over time, while useful, does not replace the fundamental need to focus on achieving augmentation. By synthesizing task-based economic theory with the concept of generalization difficulty, our work offers a scalable and low-cost methodology for organizations to forecast the economic implications of their choices. This provides a foundation for testing skill policies ex ante and for scaling experiments to study more general policy implications as candidates for future empirical validation.
Preamble
The title of this manuscript is intended to be thought-provoking, not to take a side in the debate over human versus machine superiority. It is, instead, a tribute to technology as a product of the human ability to learn, invent, and progress. While machines can outperform humans in countless tasks, they can also unpredictably fail, with consequential erosion of economic utility. At the same time, human performance may be less prone to surprise and error, but less capable to capture complex patterns and to achieve the economic benefits of high execution frequency and scale.
When a predictive model is created, whether deployed by the human mind or through a machine, its past performance serves as an estimate of future quality, assuming future conditions resemble the past. However, the real world is inherently unpredictable and doesn’t always replicate historical data. Consequently, the true value of a skill is revealed by its ability to handle the inevitable surprises of the real world. This presents a core decision-making problem for organizations of all kinds: is the best path to go all-in with automation, stick to traditional human-only methods, or find a way to blend both human and artificial skills?
This work takes inspiration from this simple question to walk a thin line, building upon a robust optimism in machines, in the hope that we (the humans) can learn how to use them in the best way. Naturally, as defining what is best goes beyond our abilities and scope, we use the principles of economics to determine measures of performance, output, value, error and utility as a first step which does not exclude to adopt other metrics.
1 Introduction
Prominent research analyzing the impact of technology over skill demand shows that computer capital allocations tend to result in i) automating well-defined and rule-based routine tasks, ii) complementing human skills in tasks requiring abstraction, problem-solving, or creativity [Autor et al., 2003], and iii) supplementing human judgment [Autor, 2024], with interesting experimental evidence in [Peng et al., 2023], [Noy and Zhang, 2023], [Brynjolfsson et al., 2025], provided the ability of using artificial capabilities effectively [Agarwal et al., 2023] [Autor, 2024].
At the same time, technological advancements in statistical learning methods, combined with increasingly affordable computational capabilities, have allowed artificial skills to achieve remarkable performance in tasks where automation through traditional rule-based approaches have historically struggled. Notable examples include image recognition and computer vision [Krizhevsky et al., 2012], natural language processing [Vaswani et al., 2017], game play [Silver et al., 2017], skin cancer diagnostics [Esteva et al., 2017], recommendation systems [He et al., 2017] and protein-folding research [Jumper et al., 2021], just to mention some of the most widely recognized examples. Moreover, language models deployed through deep neural networks [Brown et al., 2020] [Hoffmann et al., 2022] [Team et al., 2023] [Touvron et al., 2023] have contributed to amplify expectations regarding the present and future performance of artificial skills in the execution of a wide variety of cognitive tasks, making them extremely accessible to a broad audience through user-friendly interfaces. These examples emphasize the potential of technological progress to expand the set of tasks that can be automated [Acemoglu, 2025], even beyond the simpler and more compartmentalized use cases.
However, while progress in machine skills relentlessly moves forward, there seems to be tasks were machines still struggle to achieve the robustness and generality of human skills (see, for example, [Lewis and Mitchell, 2024], [Mirzadeh et al., 2024], and [Wu et al., 2024]), and the questions whether and when machines will crash the barrier of meaning [Rota, 1985] [Mitchell, 2019] still seems to remain open. Undoubtedly, machines already outperform humans in many tasks, which can surely be automated with net positive economic and societal impact under appropriate guidance, but many other situations in everyday life require levels of abstraction, analogy-making and concept slippage [Mitchell, 1993] that make the quest for full delegation to artificial skills still challenging.
Under the constantly evolving yet still persisting gap between human and artificial skills, a compelling alternative approach is to create a synergy, combining the intuitive and abstract reasoning of the human mind with the typical computational power of machines. The idea of humans and machines working in concert to achieve more than either could alone is not a new frontier, but a well-established concept that has inspired artists and scientists for generations.
As early as the late 1940s, artist Vera Molnar pioneered a collaborative approach with computers to enhance her creative process. In her conception, computational skills represent an opportunity to explore the extensive space of creations emerging from a set of pre-determined constraints, whose search exceeds human possibilities, until interesting and surprising outputs crystallize the interest of the artist. This conversational method [Baby et al., 2021], as she called it, resonates powerfully with the modern era of AI, as today’s innovators face a similar challenge to Molnar’s when striving to combine humans and machines with augmenting effects.
The concept of augmentation in human-machine interaction has been historically characterized in terms of amplification [Ashby, 1956], collaboration [Licklider, 1960] and integration [Engelbart, 1962]. Although these seminal characterizations substantially differ in the way human and machine skills contribute to the ensemble, both interpret the idea of augmentation in terms of collaborative partnership, or integrated system, aimed at achieving more than the sum of its individual parts.
More in details, augmentation can be segmented in four categories (sensory, physical, cognitive, and social) and three dimensions (replicating ability, supplementing ability, and exceeding ability) [De Boeck and Vaes, 2024]. In this context, replication pertains the ability of the ensemble to provide or restore a natural human capability, while supplement refers to an improvement or extension of a natural human capability. In contrast, an exceeding ability characterizes the achievement of a performance that is not naturally within reach for the human alone. By instance, glasses represent a case of replication, while low-magnification binoculars are a good example of supplement for human vision, whereas electron microscopes allow the human eye to exceed natural capabilities.
In a more recent treatment [Appel et al., 2025], augmentation is also analyzed through the perspective of i) learning, when the aim of human-machine interaction is to obtain information or explanations about the topic of interest, ii) task iteration, when users collaboratively partner in the execution of the task, and iii) validation, when the interaction is meant to provide a feedback to the human. These augmentation perspectives contrast with automation, either in a so-called directive mode (the machine completes the task with minimal interaction with the human agent), or by means of feedback loops (humans to support the machine in executing the task by providing feedback, when needed).
As we see, the problem of identifying and acquire (or create) the necessary skills to execute a task may not completely resolve the problem of maximizing economic results with a limited amount of resources available. How do machines handle surprising situations? Do their errors outweigh the benefits of automation? Is augmentation important when deploying a human-machine strategy, or is it better to focus on cost effectiveness? Answering these complex questions requires a rational approach that considers how performance is affected by a blurred spectrum of uncertainty. Also, while we may rapidly take a qualitative position with respect to these questions, quantitatively measuring the economic effects of human-machine interaction is a much more complex challenge, and the specific topic we explore with our work.
In this context, from an organizational perspective, and despite of many outstanding achievements, the administrative behavior [Simon, 1947] required to determine optimal capital and labor allocations under uncertainty seems far from straightforward, making skill policy decisions very complex. This appears particularly interesting when focusing on cognitive tasks, because compartmentalizing them is harder than manual or routine ones. Then, with less control over task content, more room is left for additional uncertainty at inference time with respect to training and test, thus making performance diverge from expectations, and making options of capital allocation more risky. In contrast, the option of generically favoring labor allocations under assumption of their superior ability of facing surprising situations simply leaves the economical benefits of automation behind, thus making it an equally ineffective one-fits-all solution for the problem. Inevitably, pondering these two options towards an optimal choice would require to know how uncertainty will unfold, which falls outside the ability of any decision maker in a real-world setting. This is the specific problem we present to the reader, at that we want to analyze through our work.
Luckily, we benefit from an impressive body of foundational research describing the interplay of technology and skills in a task-based framework [Autor et al., 2003] [Acemoglu and Autor, 2011], the impact of technological change on skills and labor [Acemoglu and Restrepo, 2018] [Acemoglu and Restrepo, 2019], and the dynamics behind the learning of technological solutions from data as seen through the lens of Economics [Farboodi and Veldkamp, 2021]. Moreover, we ground our understanding of the opportunities, challenges and methods of statistical learning as illustrated in [Breiman, 2001b], [Hastie et al., 2009] and [James et al., 2013], with special attention to ensemble learning [Dietterich, 2000] [Dietterich et al., 2002] and error decomposition, which may represent essential element’s in the toolkit of the modern decision-maker. Also, we make substantial reference to the approach and language used in [Chollet, 2019] when it comes to interpreting the relationship between performance and generalization difficulty, connecting concepts originating in the field of algorithmic information theory and computer science with seminal ideas traditionally belonging to the theory of decision-making under uncertainty and bounded rationality ([Simon, 1955], [Kahneman, 2011], [Kahneman and Tversky, 2013] and [Kahneman et al., 2021]). This is the complex and faceted canvas on which our work sketches some thin line, which we encourage the reader to see critically.
2 Problem framing
Our objective is to evaluate the effectiveness of different skill policies on task execution, under the effect of unpredictable task difficulty. To this end, we assume that task performance is not a deterministic outcome, but an emergent property shaped by complex, stochastic interactions between an agent’s skill, the nature of the task, and the challenge of generalizing learned skills to new problems. These dynamics are difficult to capture with traditional, closed-form mathematical equations.
To address this, we adopt a Monte Carlo simulation approach. This method allows us to model the process of task execution from the bottom-up, under various endogenous conditions, leveraging stochastic effects in the design of the interaction between skills and tasks content. Through simulations, we want to observe the organization through the lens of the distributions of task-execution performance, production output, production value and utility as measures of bottom-line economic impact accounting for the cost of error modeled as a tail risk.
Then, we take two distinct analytical perspectives, moving from specific to general. First, we leverage the model’s capabilities to examine the potential impact of the alternative skill policy options in the light of specific setup conditions. In this use case, the model represents a tool for decision-makers in need to explore the relatively small search space of their interest.
Second, we expand our focus to a more extensive experiment, enabling the analysis of a larger set of data representing a variety of conditions, to identify more general insights. To this second end, we make also use of meta-modeling, to cross-check the importance of our model’s features towards predicting the performance and bottom-line economic utility of skill policy options. In this second perspective, the model aims to be a scalable and relatively low-cost tool to explore vast search spaces to distill general recommendations or policy implications.
We are aware that an in-silico approach can have limitations. At the same time, we conjecture that some intricate and complex interactions may be hardly tractable through the lens of top-down modeling. Therefore, we hope that our effort can serve as a useful complement to other treatments. With this in mind, we encourage feedback and comments from the reader. To this end, our source code is available with a MIT license upon request.
3 Referential frameworks
The present work builds upon the model proposed in [Farboodi and Veldkamp, 2021], which defines the quality-adjusted production output of a firm $i$ at time $t$ as
$y_{i,t}=A_{i,t}k_{i,t}^{\alpha},$ (1)
where $k_{i,t}^{\alpha}$ represents the units of output of quality $A_{i,t}$ produced by leveraging $k_{i,t}$ units of capital, and $A_{i,t}=g\big((a_{i,t}-a_{i,t}^{opt})^{2}\big)$ is function of the squared distance between $a_{i,t}$ (the production technique of choice for firm $i$ at time $t$) and the optimal technique $a_{i,t}^{opt}=\theta_{t}+\epsilon_{a,i,t}$}, with $\theta_{t}$ and $\epsilon_{a,i,t}$ being the jointly-observed persistent and transitionary components of the optimal technique, and $g$ a strictly decreasing function.
This formulation, by means of the introduction of the squared distance between the actual and the optimal production skill, makes the role of technological performance explicit in the determination of production output of automated tasks. Also, this framework offers the opportunity to explore a perspective that may complement the treatments illustrated in [Autor and Thompson, 2025], where a worker is able to execute a tasks only if the worker’s expertise level is equal or higher than the one required by the task (the hierarchical expertise assumption), and automation deploys below a certain threshold of task complexity (the hierarchical automation assumption, a condition also illustrated in [Acemoglu, 2025] as effect of the comparative advantage offered by labor with respect to capital in newer or complex tasks).
Our work is centered on the idea that performance emerges from the interaction between task content and skills, either human or artificial, and that part of the task difficulty may depend on the level of generalization required. To model these concepts, we draw heavily from the concepts illustrated in [Chollet, 2019], with particular regard to generalization difficulty, i.e. the increased complexity introduced by additional levels of uncertainty presenting at execution time with respect to learning time. The foundational concepts of Chollet’s framework serves as the second theoretical pillar of our work.
4 Model
To the aim of our work, we elaborate the framework illustrated in Section 3 with additional treatments, which we synthetically illustrate in the following sections. Separately, we summarize the process of obtaining the technological solution $a_{i,t}$ in terms of a decision-making problem whose candidate solutions are denoted as $Sol_{i,t,c,j,[,d,q]}$, with a notational convention aligning with [Chollet, 2019] -see Appendix, Section A.4.
4.1 Loss ($\mathcal{L}$) and performance ($\theta$)
First, we move by abstraction to defining the distance between the actual and optimal skills in terms of loss ($\mathcal{L}$), such that the quality $A_{i,t}$ can be expressed as
$A_{i,t}=g\big(\mathcal{L}_{i,t}\big),$ (2)
where $\mathcal{L}_{i,t}=f(a_{i,t},a_{i,t}^{opt})$ and $f$ is any appropriate function measuring the absolute gap between the actual and optimal technological skills. By doing this, we disentangle the model from the specific choice of square distance, opening the possibility to consider alternative options including measures of divergence or dissimilarity (also, if necessary, multiple measures of loss may be combined). Then, when the range of potential values of $\mathcal{L}_{i,t}$ is known, we obtain a normalized loss $\mathcal{L}^{\prime}_{i,t}\in[0,1]$ through the normalization process of choice, by instance
$\mathcal{L}^{\prime}_{i,t}=\frac{\mathcal{L}_{i,t}-\min(\mathcal{L}_{i,t})}{\max(\mathcal{L}_{i,t})-\min(\mathcal{L}_{i,t})}.$ (3)
Next, we define performance the skill level [Chollet, 2019] achieved by firm $i$ at time $t$ as
$\theta_{i,t}=(1-\mathcal{L}^{\prime}_{i,t}),$ (4)
and
$y_{i,t}=g^{\prime}\big(\theta_{i,t}\big),$ (5)
where $g^{\prime}:\theta_{i,t}\in[0,1]\rightarrow A_{i,t}{kl}^{\alpha}_{i,t}\in[0,1]$ is a monotone and non decreasing function mapping performance to quality-adjusted output (the combined evaluation of $A_{i,t}$ and ${kl}^{\alpha}_{i,t}$ through the function $g^{\prime}$ is a desired behavior of our model, to streamline the formal relationship between performance and quality-adjusted output).
As a result, for a generic task of interest at time $t$, the relationship between the distribution of values observed for the quality-adjusted task output and the distribution of the performance achieved in the execution of that task by the firm $i$ deploying the technological solution of choice $a$ are modeled by the function $g^{\prime}$. Then, the problem of estimating the distribution of quality-adjusted production output $y$ can be tackled by modeling the distribution of $\theta$ and the form of $g^{\prime}$. In the following sections we propose how to do this.
4.2 From performance to quality-adjusted output
Our modeling strategy is based on the principle that, given the technological solution of choice (a skill, or a predictive model), the level of performance assessed at test time may differ from the level of performance measured at inference time. The causes of this phenomenon can be many, but we specifically focus on two of them: first, the skill may be exposed to situational instances introducing new and additional levels of uncertainty with respect to what experienced at training time, thus reducing the model ability to respond accurately; second, when more than one model is available, the decision regarding which model to deploy, individually or in an ensemble, can condition the final prediction.
We briefly describe our treatment of these two conditions in the following Sections 4.2.1 and 4.2.2. Then, in Section 4.2.3 we describe the modeling strategy adopted to jointly implement them in the measure of skill performance and quality-adjusted task output.
4.2.1 Skill policy ($c$)
In the present work, we focus our attention on productions that can be obtained through cognitive tasks exclusively (we leave open the opportunity to relax this assumption in further steps), and we extend the definition of production technique to include models operated by the human mind.
In this frame of reference, and under the assumption of labor and capital being perfect substitutes while potentially providing different comparative advantages [Acemoglu and Autor, 2011] [Acemoglu and Restrepo, 2018] [Acemoglu, 2025], we specifically define skill policy the choice made by each firm with regards to the combination of labor and capital allocations to the task. In a first instance, we consider three distinct skill policies, similarly to what illustrated in [Agarwal et al., 2023]:
  1. 1.
    Skill policy $H$ (human-exclusive): the firm’s choice to execute the production task of interest by exclusively leveraging labor allocations (no use of technology altogether), resulting in the quality-adjusted output $y_{i,t,l}=g^{\prime}\big(\theta_{i,t,l}\big)$}, where $\theta_{i,t,l}$ denotes the skill level achieved by uniquely leveraging labor and models deployed by the human mind;
  2. 2.
    Skill policy $M$ (machine-exclusive): the firm’s choice to execute the production task of interest by exclusively leveraging capital allocations (automation), resulting in $y_{i,t,k}=g^{\prime}\big(\theta_{i,t,k}\big)$}, where $\theta_{i,t,k}$ denotes the skill level achieved by uniquely allocating capital;
  3. 3.
    Skill policy $HM$ (human-machine): the firm’s choice to combine labor and capital, which determines a variety of ways in which human expertise and technological capabilities can interact, potentially leading to augmentation, and resulting in a quality-adjusted output computed as
    $y_{i,t,kl}=g^{\prime}\big(a(\theta_{i,t,k},\theta_{i,t,l},\gamma_{kl})\big),$ (6)
    where $a$ is the function computing the performance of the hybrid allocation of capital and labor deployed as an ensemble and $\gamma_{kl}$ represents a factor, or a set of factors, influencing the augmenting effects of the function $a$ (at the core of the skill policy $HM$ is the idea that the performance of the ensemble can surpass the performance of each contributing model when individually considered, which is a basic definition for augmentation); prominent examples of ensemble learning techniques include Bootstrap Aggregating (Bagging) [Breiman, 1996], Random Forest [Breiman, 2001a], and AdaBoost [Freund et al., 1996]. From an economic perspective, the role of the function $a$ can be interpreted as a term modeling interaction or delegation policies.
Synthetically, for a firm $i$ deploying a skill policy $c$ at execution epoch $e$, we denote as $y_{i,c,e}$ the quality-adjusted output of production resulting from the achievement of the skill level (performance) $\theta_{i,c,e}$.
4.2.2 Task difficulty ($d$)
In our model architecture, we define task difficulty as a task property revealing itself through instances characterized by different levels of uncertainty faced at inference time $t+e$ ($e>0$) with respect to training/testing time $t$.
The effort required to the skill of interest to keep the same skill level (performance) when confronting with increasing uncertainty with respect to training time is typically referred to as generalization difficulty, which can be distinguished in local generalization and broad generalization [Chollet, 2019]. In a first instance, we denote difficulty levels as $d\in D=\{Low,Med,High\}$, assuming that each level requires a specific ability of broad generalization.
4.2.3 Performance as a probability distribution
To jointly account for the effects of task difficulty and skill policy, we represent performance in terms of a set of probability distributions $P(\theta)_{c,d,e}$ (see Figure 1), which are specific for each skill policy $c$ confronting with a task instance of difficulty $d$ at epoch $e$. Additionally, we model these distributions in a way that an evolution over epochs is allowed, to represent that skills can change (for example: while a machine skill may be static over epochs, until a new training is done, human skills constantly evolve; this is an interesting behavior to study, requiring appropriate modeling).
In a first modeling instance we consider the $P(\theta)_{c,d,e}$ distributions as universal, thus shared by all firms (however, this constraints can be relaxed in future steps of analysis). Then, the performance achieved by the firm $i$ adopting skill policy $c$ at epoch $e$ in the execution of a task instance presenting difficulty $d$, denoted as $\theta_{i,c,d,e}$, is determined by random draw based on the probability distribution $P(\theta)_{c,d,e}$, with a resulting quality-adjusted output $y_{i,c,d,e}=g^{\prime}(\theta_{i,c,d,e})$.
Figure 1: $P(\theta)_{c,d,e}$ for for all skill policies and difficulty levels at epoch $e=0$ and e=E1e=E-1. The probability densities presented in this example here are modeled through Beta distributions to describe an evolution of the skill $H$ over epochs, while the skill $M$ is static over time.
4.3 Margin of contribution and value
Moving from quality-adjusted units of output to value requires to consider the economic elements at play. As a deep exploration of this topic goes beyond the scope of our work, we synthetically leverage the net margin of contribution that the firm obtains from each unit of quality-adjusted output (denoted as $mc_{c}=\{mc_{H},mc_{M},mc_{HM}\}$, with $mc_{c}\in[0,1],\forall c$), which is specific for the skill policy of choice.
While imposing no constraints to the function used to compute $mc_{c}$, we assume it to depend on i) the price recognized by the market for a unit of production with quality $A_{i,t}$, and on ii) the expenditures related to capital deployments alongside recurrent labor costs associated to the skill policy of choice and, consequently, to the technological solutions at work, allocated over their economic life to each unit of productive output according to the preferred criteria. As a result, we define the value of quality-adjusted production output, denoted as $v$, as
$v_{i,c,d,e}=max(0,min(1,y_{i,c,d,e}*mc_{c}*\delta_{c,e})),$ (7)
which is clipped to the range [0,1].
As the reader will promptly notice, the parameter $\delta_{c,e}$ depends from both $c$ and $e$ to denote the potential evolution of the margin of contribution through epochs, selectively for each skill policy. In a first instance, we compute $\delta_{c,e}$ by linear interpolation as
$\delta_{c,e}=interpolate(1,1+\delta_{c},e,E),$ (8)
where $\delta_{c}=\{\delta_{H},\delta_{HM},\delta_{M}\}$ are configuration parameters determining the increase or decrease of the margin of contribution over epochs, modeling the progress of the firm in walking the skill learning curve, or the firm’s ability to improve the cost effectiveness of skill development and operations, or the evolution of the market price for the task production under consideration ($\delta_{c}>0$ determines an increase, i.e. a unit of quality-adjusted production increases through epochs, while $\delta_{c}<0$ determines a decrease).
It is worth noticing that the generality of the $\delta_{c}$ parameters comes at the price of a simplification, because no specific assumption about the factors influencing the organization efficiency (or productivity) is made. For example, as illustrated in [Farboodi and Veldkamp, 2021], learning production techniques from data is one path that organizations can walk to increase their efficiency, and with data being a by-product of the organization’s activity, the more productive the firm is, the more data it can generate and store, thus entering the so-called data feedback loop. Naturally, modeling this effect goes far beyond setting the slope of a linear growth. The design we propose through the $\delta_{c}$ parameters represents a first step, while a non-linear function of cumulative data or experience may be a promising direction for future developments.
4.4 Error and net utility
Additionally, we introduce the element of error as a shock conditioned on performance, implying that the economic impact of a low effectiveness is not only measured in terms of reduced output, but also in terms of an economic penalty triggered beyond a minimum performance threshold $t_{err}\in[0,1]$. This threshold can be interpreted in terms of quality control threshold or level of safety tolerance.
In combination with $t_{err}$ we introduce the parameter $c_{err}\in[0,1]$, which defines the cost of error in terms of fraction of input units ($c_{err}$ may represent a variety of phenomena, including the cost of the necessary corrections, the cost of repair for eventual damages, a financial penalty, a reputation damage, or the cost of a product recall). Combining these factors, we compute
$err_{i,c,d,e}=\begin{cases}(1-\theta_{i,c,d,e})c_{err}&\text{if }(1-\theta_{i,c,d,e})\geq t_{err},\\ 0&\text{otherwise.}\end{cases}$ (9)
Finally, utility, denoted as $u$, is computed as
$u_{i,c,d,e}=v_{i,c,d,e}-err_{i,c,d,e}.$ (10)
5 Simulation model
By leveraging the properties of the model illustrated in the previous Section 4, we deploy the method of Monte Carlo simulation to produce synthetic data. The scope of this section is two-folded: in a first part, we illustrate the logics governing the simulation, and we formalize the description of the simulation output serving as input for the subsequent analyses; in a second part, we discuss the challenges underlying the use of simulation as a tool for observing emerging phenomena, detailing the reasons why we think that this approach can be viable and valuable for our research.
5.1 Underlying principles and methodology
Without sacrificing generality, we consider a single production task being executed by a number of firms involved over multiple epochs. The parameters given as input for the simulation algorithm are here briefly discussed.
5.1.1 Size of a simulation
We define the size of the simulation by means of three parameters:
    NN, the number of firms, individually denoted by the index $i=0,\ldots,N-1$;
    $E$, the number of epochs, individually denoted by the index $e=0,\ldots,E-1$;
    $K$, the number of runs for which each simulation is repeated (replications), individually denoted by the index $k=1,\ldots,K$.
Given a computational environment and its related capabilities, the size of the simulation determines the time and cost required to execute it and generate the corresponding data and eventual visualizations.
5.1.2 Probability distributions for skill policies and task difficulties
At simulation start, every firm is assigned a static skill policy $c\in\{H,HM,M\}$, according to the probability distribution $P_{c}=(P_{c=H},P_{c=HM},P_{c=M})$. At each epoch, all the firms are assigned a task, whose difficulty $d\in\{Low,Med,High\}$ is determined by random draw, according to the distribution of probabilities $P_{d}=(P_{d=Low},P_{d=Med},P_{d=High})$.
5.1.3 Probability distributions for task-execution performance
Given a firm deploying skill policy $c\in\{H,M\}$ and a task with difficulty $d$, the skill level $\theta$ achieved in the execution of the task at epoch $e$ is determined by random draw, according to the distribution $P(\theta)_{c,d,e}\sim Beta(\alpha_{c,d,e},\beta_{c,d,e})$, with $c\in\{H,M\}$. In our first modeling instance, we chose the Beta distribution to benefit from the support in $[0,1]$ and from the flexibility in approximating a wide variety of distributions, including uniform, normal, right-skewed or left-skewed ones; however, to approximate the real-world case of need, any kind of distribution may be theoretically applicable, provided appropriate clipping to $[0,1]$.
$\forall c\in\{H,M\}$ [^1] We use this notation for simplicity because $c$ can assume only two values differing from $HM$. However, in the more general case of more then three skill policies, the notation $\forall c\in\{H,M\}$ should be replaced by $\forall c\neq HM$., through their parameters $\alpha_{c,d,e}$ and $\beta_{c,d,e}$, the distributions $P(\theta)_{c,d,e}$ can be dynamic over epochs. The point value of $\alpha_{c,d,e}$ and $\beta_{c,d,e}$, $\forall c\in\{H,M\}$ is determined, $\forall e\notin\{0,E-1\}$, by interpolation with respect to initial and final values given as input, according to the Algorithm 3 illustrated in Appendix, Section A.1. We generically denote the values to be given as input as pairs $(\alpha,\beta)_{c,d,e}$, $\forall$ combination of $c\in\{H,M\}$, $d\in\{Low,Med,High\}$ and $e\in\{0,E-1\}$.
5.1.4 Performance for the human-machine skill policy (function $a$)
With the performance for skill policies $H$ and $M$ being determined by random draw from $P(\theta)_{c,d,e}$, the performance for the human-machine skill policy is computed according to the following options for the function $a:(\theta_{c=H},\theta_{c=M},\gamma_{HM})\rightarrow\theta_{c=HM}$, $\forall i,d,e,k$:
    $a=mean$, representing the case of a firm that cannot predict which skill (human or artificial) will perform better in the specific task instance, i.e. ($\forall i,d,e,k$):
    $\theta_{c=HM}\xleftarrow{}mean(\theta_{c=H},\theta_{c=M});$ (11)
    $a=max$, an optimistic (upper) bound for $a=mean$, representing the case of an omniscient firm which is able to optimally delegate task execution to either human or artificial skills, i.e. ($\forall i,d,e,k$):
    $\theta_{c=HM}\xleftarrow{}max(\theta_{c=H},\theta_{c=M});$ (12)
    $a=min$, a pessimistic (lower) bound for $a=mean$, representing the case of a firm systematically opting for the implementation of the solution achieving the lower performance, i.e. ($\forall i,d,e,k$):
    $\theta_{c=HM}\xleftarrow{}min(\theta_{c=H},\theta_{c=M});$ (13)
    $a=collaborate$, representing the case of human and artificial skills to interact towards potential augmentation, computing as ($\forall i,d,e,k$):
    $\theta_{c=HM}\xleftarrow{}\min(1,mean(\theta_{c=H},\theta_{c=M})*\gamma_{HM}));$ (14)
    where $\gamma_{HM}$ is the augmenting factor applied, with a cap for the final augmented performance set to 1;
    $a=superpower$, representing the case of strong augmentation, computing as ($\forall i,d,e,k$):
    $\theta_{c=HM}\xleftarrow{}\min(1,max(\theta_{c=H},\theta_{c=M})*\gamma_{HM})).$ (15)
Clearly, the practical realization of human-machine interaction in a superpowersuperpower-like way is the Holy Grail of many AI initiatives: a process which is designed not only to route a task to the most effective agent, but the realization of a synergy that surpasses the sum of the individual contributions. This represents a proxy for a sophisticated and real-time skill-orchestration. Unfortunately, describing how deployments like collaboratecollaborate or superpowersuperpower can be realized in practice is outside the scope of our work, and probably beyond the authors’ abilities, also because their design inevitably requires a context-specific approach.
5.1.5 Quality-adjusted output (function $g^{\prime}$)
In a first modeling iteration, for the function $g^{\prime}:\theta\rightarrow y$ mapping skill level to quality-adjusted production output, we consider the following options and their respective parameters, with $y$ clipped to the range $[0,1]$ (see Figure 2):
    linear: $y=slope\text{ }\theta+intercept$;
    logistic: $y=1/(1+e^{-k(2\theta-1)})$, penalizing lower performance with respect to a linear relationship;
    inverse_logistic: $y=[k-\ln(1-\theta)-\ln(\theta)]/(2k)$, penalizing higher performance with respect to a linear relationship;
    exponential: $y=(base^{\theta}-1)/(base-1)$, with $base\neq 1$;
    power: $y=\theta^{exponent}$;
    logarithmic: $y=\ln{[1+\theta*(log\_base-1]})/\ln(log\_base)$;
Figure 2: Instances of the function $g^{\prime}$, with some exemplary parameters
This brief list represent a relatively flexible set of option to fit a variety of common situations. More in general, beyond our arbitrary proposals, the decision-maker may introduce the preferred function of choice based on field knowledge. The function providing with the most accurate estimation of the real-world phenomena under analysis should be preferred in the design on the simulation algorithm.
5.1.6 Margin of contribution
Once the quality-adjusted output $y_{i,c,d,e,k}$ is determined (units, normalized to [0,1]), the value $v_{i,c,d,e,k}$ is obtained by applying the equation (7) $\forall k$ based on the margins of contribution $mc_{c}=\{mc_{H},mc_{HM},mc_{M}\}$, with $mc_{c}\in[0,1]$ being provided as input parameters. Also, the parameters $\delta_{c}=\{\delta_{H},\delta_{HM},\delta_{M}\}$ are provided as input, determining the increase, or decrease, of the margin of contribution over epochs.
5.1.7 Cost of errors
To determine $err_{i,c,d,e,k}$ according to Equation (9), the parameters $t_{err}\in[0,1]$ and $c_{err}\in[0,1]$ are required. The performance threshold $t_{err}$ triggers the computation of the cost of error when $\theta_{i,c,d,e,k}<t_{err}$, and the actual unitary cost $c_{err}$ determines the size of the consequent economic penalty incurred by the firm.
5.1.8 Simulation algorithm
Accordingly, $\forall k$ (see Appendix, Section A.1, Algorithm 4), if $e=0$ each firm is assigned a skill policy (static through epochs), according to the probability distribution $P_{c}$. Then, $\forall e\in[0,E-1]$:
    Each firm is assigned a task with difficulty level $d$, according to the probability distribution $P_{d}$ (the probability is static through epochs, and draws occurring at each epoch are independent);
    Each firm executes the production task with difficulty $d$, and the resulting skill level $\theta_{i,c,d,e,k}$ is determined according to the provisions illustrated in Section 5.1.3 and 5.1.4, given the firm’s skill policy $c$;
    Then, each skill level $\theta_{i,c,d,e,k}$ is passed to the function $g^{\prime}$ to compute the quality-adjusted output $y_{i,c,d,e,k}$;
    Next, each quality-adjusted output $y_{i,c,d,e,k}$ is converted in an economic value $v_{i,c,d,e,k}$ by means of the appropriate margin of contribution $mc_{c}$;
    Also, a cost of error $err_{i,c,d,e,k}$ is computed;
    Next, the utility $u_{i,c,d,e,k}$ is obtained as $v_{i,c,d,e,k}-err_{i,c,d,e,k}$, $\forall i,e,k$;
5.1.9 Simulation output
As a result of each simulation $s$, the following sets of values are collected:
    $\Theta^{s}=\{\theta_{i,c,d,e,k}^{s}\}$, measuring the performance, or skill level;
    $Y^{s}=\{y_{i,c,d,e,k}^{s}\}$, quantifying quality-adjusted output;
    $V^{s}=\{v_{i,c,d,e,k}^{s}\}$, determining the economic value of quality-adjusted output;
    $ERR^{s}=\{err_{i,c,d,e,k}^{s}\}$, computing the cost of error;
    $U^{s}=\{u_{i,c,d,e,k}^{s}\}$, recording the net utility resulting from task execution.
In this context, for notational convenience, we generically refer to these sets as $\Lambda^{s}=\{\lambda_{i,c,d,e,k}^{s}\}$, with $\lambda\in\{\theta,y,v,err,u\}$ and, consequently, $\Lambda\in\{\Theta,Y,V,ERR,U\}$.
The output resulting from a single simulation may be analyzed individually and the distributions of the values in the sets $\Lambda$ visualized (one example is presented in Figure 4). Naturally, within an individual simulation, the index $s$ in the notations can be dropped for simplicity and ease of understanding.
Figure 3: Schematic representation of an experiment combining multiple simulations with different input parameters, and the subsetting/processing of the resulting data.
Figure 4: Example of visualization of the result of an individual simulation. For every sector of the grid, a distribution of values is visualized as a box plot (values grouped by skill policy $c$, with $d$ as hue). In rows, from top to bottom, we have the distributions of the values in the sets $\Theta$, $Y$, $V$, ERRERR and $U$. In columns, from left to right, we have i) the distribution of values in the first epoch, ii) the distribution of values in the last epoch, iii) the distribution of all values, and iv) the arithmetic means of the values by skill policy and task difficulty (under the title of each plot in the last column if the arithmetic mean of all the values, i.e. $\overline{\Lambda^{s}},\forall\Lambda\in\{\Theta,Y,V,ERR,U\}$
Additionally, we define:
    $\Lambda^{s}_{c_{0}}=\{\lambda_{i,c,d,e,k}^{s}:c=c_{0},\forall i,e,d,k\}\subseteq\Lambda^{s}$, i.e. the subsets of values characterized by the same skill policy $c_{0}$;
    $\Lambda^{s}_{c_{0},d_{0}}=\{\lambda_{i,c,d,e,k}^{s}:c=c_{0}\land d=d_{0},\forall i,e,k\}\subseteq\Lambda^{s}_{c_{0}}$, i.e. the subsets of values characterized by the same skill policy / task difficulty pair $(c_{0},d_{0})$;
and
    $\mu(\Lambda^{s})$, or $\overline{\Lambda^{s}}$, is the arithmetic mean of the values in the set $\Lambda^{s}$;
    $\sigma(\Lambda)$ is the standard deviation of the values in the set $\Lambda^{s}$;
    $\rho(\Lambda^{s})=\max(\Lambda^{s})-\min(\Lambda^{s})$ is the range of the values in the set $\Lambda^{s}$;
    $IQR(\Lambda^{s})$ is the interquartile range (IQR) of the values in the set $\Lambda^{s}$;
    $SK(\Lambda^{s})$ is the skewness of the probability distribution of the values in the set $\Lambda^{s}$,
where $\mu$, $\sigma$, $\rho$, $IQR$ and $SK$ are also generically denoted as summary statistics $\omega$, with the same logic being applicable also to the sets $\Lambda^{s}_{c_{0}}$ and $\Lambda^{s}_{c_{0},d_{0}}$. Then, the distributions of $\omega(\Lambda^{s})$, $\omega(\Lambda_{c}^{s})$ and $\omega(\Lambda_{c,d}^{s})$ can be visualized by means of box plots. A schematic representation of the process is illustrated in Figure 3, while Figure 5 provides an example of visualization of an experimental output.
Figure 5: Example of visualization of the result of an experiment consisting of multiple simulations. Our typical visualization features a variable number of rows, depending on the desired distributions of interest, but always three columns. Rows represent the choice of $\omega(\Lambda^{s})$ (in this example: $\mu(\Theta^{s})$, $\sigma(U^{s})$, $\rho(Y^{s})$, $IQR(U^{s})$, $SK(\Theta^{s})$). Columns, instead, represent the progressive break down of values in $\Lambda^{s}$: in the first column on the left, we visualize the entire distribution $\omega(\Lambda^{s})$, in the second we have $\omega(\Lambda_{c}^{s})$ and in the third we have $\omega(\Lambda_{c,d}^{s})$. Each visualization represents a combination of overlapping violin plot and box plot.
5.2 Rationale for the Monte Carlo method
A fundamental critique of computational simulation is that its results risk to be a mere consequence of the hypotheses and assumptions encoded in the model (the tautology critique), unless reasons justify its use. We believe that acknowledging this risk is a necessary step, together with a detailed rationale, to support the choice of simulation in the use case of interest.
To this end, we articulate two main arguments: i) the capacity of simulation to generate non-trivial emergent phenomena which cannot be easily deduced from the top down, and ii) the flexibility and ease of implementation of a wide range of experimental settings, which can have positive consequences on the ability to explore the phenomena of interest. These, through the appropriate methodological safeguards of verification and validation, can allow to ground insights in objective reality, thus contributing to strengthen the validity of the findings and avoid the fall in the shadow of tautology machines.
The most direct counter to the tautology critique is the concept of emergence. This describes a phenomenon where novel, macro-level patterns and behaviors arise from the simple, local interactions of individual components. These higher-level properties are not explicitly programmed into the system, but an unexpected consequence of more complex behavior of the parts.
To formalize this, a very useful treatment is illustrated in [Chalmers, 2006], with the distinction between weak emergence and strong emergence. Both these concepts describe the rise of high-level phenomena from low-level dynamics, but they differ by their deductibility and predictability. Strong emergence denotes high-level phenomena which cannot be deducted from low-level domain principles. Weak emergence, instead, characterizes phenomena which are unexpected, not immediately or directly deductible from low level principles and initial conditions, but that can be theoretically traced back to them with a dedicated effort. In short, we define a pattern strongly emergent if it is detectable only upon the observation of high level effects produced by low level dynamics. In contrast, the pattern is weakly emergent if it is not obvious from basic principles or setup conditions, but it results a straightforward consequence of them upon in-depth examination. This characterization of weak emergence highlights its observer-relative nature, as it depends on the ability of the observer to deduct it from the top-down rather then from the observation of effects crystallizing from the bottom-up.
Seminal examples from the literature underscore this point. In the study of complex stochastic combinatorial optimization, for example, interesting solutions arise from the observation of coordinated collective behavior emerging from the simulation of agents with extremely simple basic capabilities at work [Dorigo et al., 1996]. Similarly, agent-based models of social phenomena show how simple rules for individual behavior can result in larger scale and counterintuitive social patterns like residential segregation [Schelling, 1971].
Another characteristic of emergence is that it heavily relies on simulation to manifest itself to the observer. Whether the emergent phenomena is inherently non-deductible from basic known principles (strong emergence) or just not easily deductible from them (weak emergence), the absence of simple analytical paths or shortcuts for the prediction of the final behavior from the initial conditions makes emergence a discovery based on the observation of the process running.
By executing an algorithm that is known, the simulation process does the work that is necessary to revealing an effect that is extremely challenging to qualify or quantify analytically. Then, a behavior or a property can emerge from simulation with relative simplicity, while its analytical top down computation could be impossible, or require relevant effort. In our use case, this concept is represented by the challenge of analytically determining the statistical description of the distributions of the values in the sets $\Theta_{c,d}^{s}$ and $\omega(\Theta_{c,d}^{s})$ based on the initial characterization of $P(\theta)_{c,d,e}\sim Beta(\alpha_{c,d,e},\beta_{c,d,e})$. While theoretically feasible, determining the moments of a probability distribution whose values are obtained as a function of values drawn from other distributions can result in an escalation of complexity that poses barriers to flexibility and ease of experimental setup.
Also, another realization of macro-level effects originating from (but not easily deductible from) causal low-level principles is when setup conditions produce shocks upon the occurrence of certain events. In our model, this is exemplified by the cost of errors. Although it’s intuitive to understand that the cost of error computes by multiplying the cost of an individual event by the estimated number of errors, quantifying this effect with precision under the influence of stochastic effects poses challenges.
In these cases, a synthetic measure of central tendency may be relatively easy to compute, but it may be reductive, as an analysis of the distribution of values rather than their mean or median value can give valuable information. In this context, the existence of the possibility to obtain an aggregate analytical estimation does not eliminate the option of simulation, especially when simulation offers the attractive options to rapidly increase the complexity of the model and easily produce distributions of values. Naturally, the use of simulation comes at the cost of resources (mainly time and compute) and approximation, which are in part related each other; however, both can be handled and mitigated with appropriate methods.
The tautology critique also fails to consider distinctive epistemic traits of simulation, whose benefits have been clearly represented in pioneering works [Axelrod, 1997] [Epstein, 2012]. Deduction begins with explicit axioms and uses formal logic to prove theorems. The outcome is a logical certainty. Induction begins with empirical data and seeks to discover general patterns or laws. Simulation, in contrast, operates by combining aspects of the two: it starts from explicit assumptions to generate synthetic data through computational experiments. Then, the task is to analyze simulated data to uncover surprising patterns and unanticipated consequences that were not obvious from the initial assumptions. This makes simulation a valuable tool for exploratory modeling, especially for complex and dynamic systems, such as those in social science or economics, where making detailed predictions is often impossible due to a variety of uncertain or intricate effects.
In these scenarios, the objective of a rigorous approach based on simulation is not to find a single definitive answer to a problem by directly hard coding the solution in the model, but to assist research in reasoning about the system itself, especially when empirical data is insufficient or not available at all. The rigor and value of a simulation-based approach may bee seen in how it is conducted, not in the approach itself: when an emergent behavior can be formulated as a testable hypothesis about the real world, thus subjected to further empirical testing, then it can be a valid option to pure deductive or inductive methods. For this very reason, the arguments for emergence and exploration require that a simulation is methodologically rigorous. The tautology machine critique often serves more as a proxy for a lack of trust in the model, not in the method itself. A robust defense, therefore, requires a clear and transparent framework for establishing model credibility. The gold standard for this is the process of Verification and Validation (V&V).
Verification and validation are distinct but interdependent: verification is the process of answering to the question ‘Did I build the thing right?’ [Pace, 2004], primarily related to the internal correctness and consistency of the model and its implementation. Validation, instead, pertains to the question ‘Did I build the right thing?’, concerning with the external credibility of the model, determining the degree to which the model is an accurate representation of the real world for its intended use. A rigorous V&V framework is a direct response to the tautology critique. Verification ensures that the output is not a tautology by mistake, while validation ensures the model is not a tautology by design, as it forces the modeler to link the model hypotheses to an empirically testable context.
As we have briefly illustrated, the true potential of simulation lies in its capacity to reveal emergent phenomena that are a non-obvious consequence of its initial rules, serving as a platform for computational experiments that generate new, falsifiable hypotheses, that can (and should) be subjected to verification and validation. With this in mind, we devote the conclusive part of this rationale to quickly review the specific assumptions and implementations characterizing our model, explaining why we think that a simulation-based approach is sound and beneficial.
5.2.1 Model validation
Logics and characterization of the results
Our framework determines the local performance $\theta_{i,c,d,e,k}$ for the firm $i$ deploying skill policy $c\in\{H,M\}$ in the execution of a task with difficulty $d$ at epoch $e$ (run $k$) by random sampling from global distributions $P(\theta)_{c,d,e}\sim Beta(\alpha_{c,d,e},\beta_{c,d,e})$. By construction, the parameters $\alpha_{c,d,e}$ and $\beta_{c,d,e}$ can evolve over epochs to represent how skills change over time. Then, a first level of complexity in implementing an analytical process characterizing the distribution of performance is represented by computing the moments of a distribution that results from the interaction of a variety of other distributions. This is even more complex for the performance achieved by deploying skill policy $HM$, which is obtained as a function of the performance that skill policies $H$ and $M$ would obtain on their own (see Equation (6)).
In this case, depending on the construction of the function $a$ in conjunction with the type of distribution of choice for $P(\theta)$, the analytic computation of the moments for $\Theta^{s}$, $\Theta_{c}^{s}$ and $\Theta_{c,d}^{s}$ can be hard or even impossible. Additionally, according to our model, the function $g^{\prime}:\theta\rightarrow y$ can be non-linear, thus posing another element of complexity to the analytical treatment of distribution characterization. Consequently, even when stochastic effects are limited to the performance and the other economic metrics can be deterministically computed, the difficulties of appropriately characterizing the distributions of performance poses factual limits to analytical implementations.
Finally, the possibility for margins of contribution to evolve over epochs, and the discontinue nature of the function triggering the cost of error make a comprehensive analytical treatment very complex.
Setup conditions
A critical consideration for our simulation approach is the selection of parameter ranges, as this choice can influence the simulation’s output and potentially bias findings. To mitigate this risk, we first establish the circumstances under which specific range limitations are justified by empirical realism rather than by subjective preferences.
In this context, an important choice pertains the margins of contribution. In a typical production environment, automation is expected to generate lower variable costs then manual labor, and to benefit from higher economies of scale. However, the positive effect of automation on productivity tends to be contrasted by the high investment costs, thus promoting the development of cost-friendly automation solutions [Jäger et al., 2016] to enhance the positive effect on total productivity. In a theoretical environment where the margin of contribution is normalized to a standard [0,1] range, empirical realism requires to bound the values of $mc_{H}$ and $mc_{M}$ to specific sub-ranges such that $mc_{H}<mc_{M}$, and define $mc_{HM}$ as a function of $mc_{H}$ and $mc_{M}$ (deploying both skills implies incurring in both the related costs). However, as the firm can progressively improve its ability to develop cost-effective technological investments, the gap between $mc_{HM}$ and $mc_{H}$ (and $mc_{M}$) may reduce over time. These real-world constraints, pondered by the specific characteristics of the use case under analysis, are essential to restrict the exploration space to the portions of it connected to situations that can be encountered in the real world, to enable and ease verification.
Also with regards to the ranges in which $\gamma_{HM}$ and $\delta_{c}$ can take their values, reasonable hypotheses are required. For example, while there is no methodological limitation to the numeric value a multiplicative gain can assume, it is unrealistic to think that it be extremely high (is it realistic to think that augmentation can produce a 10x performance increase?). Similarly, a growth or decrease of the margin of contribution over epochs seems reasonable and realistic if relatively contained, because explosive margin dynamics are rare.
6 Applications
The preceding sections have established a theoretical model and simulation framework to analyze the economic impact of different skill policies in production cognitive tasks. This section shifts from theory to application, demonstrating the practical utility of our methodology for decision-makers. The two distinct approaches presented illustrate how the model can be a versatile tool: first, for focused, micro-level analysis of specific organizational scenarios, and second, for broader, macro-level experimental exploration to derive more general policy insights. This duality highlights the model’s ability to provide both immediate, quantitative support for tactical decisions and a foundation for scalable, strategic forecasting.
6.1 Exploring a small space: the case of individual simulations
When faced with a specific decision-making task, organizations and policymakers may prioritize predicting the outcomes of the available options rather than analyzing the problem through a general framing. Unlike a broader research objective, this approach has a more compartmentalized scope that doesn’t require a full exploration of the entire solution space. Naturally, the prerequisite is that the situation and its environmental conditions can be characterized in terms of model variables.
For instance, an organization can draw upon empirical data to determine the average contribution margin per unit of quality-adjusted output generated by a specific task under a human-exclusive skill policy. Also, organizations can leverage their domain-specific expertise to estimate the cost of errors and establish the performance threshold at which these errors are likely to emerge. By strategically limiting the scope of configurations under consideration, our simulation model enables organizations to directly analyze the expected economic impact of alternative skill policies, and rapidly estimate their return on investment. This represents an entire class of inherently specific use cases, whose understanding may be easier by means of a practical example.
To this end, we describe the characteristics of a small set of realistic scenarios for our fictitious decision-maker, and then we use our model to analyze the potential quantitative economic outcomes that may arise from the choice of the skill policies of interest. Given by construction a limited number of situations to explore, we set the number of Monte Carlo repetitions to $K=10,000$ to enhance the robustness of our simulation results.
The experimental setup we propose is based on uniform probability distributions for skill policies and task difficulties, and on Beta-distributed probabilities $P(\theta)_{c,d,e}$, parametrized as illustrated in Figure 8. Next, we posit that $g^{\prime}:\theta\rightarrow y$ is a logistic function with a steepness parameter of $k=5$. Also, we assume a per-unit contribution margin of $mc_{H}=0.5$ for output produced through human-exclusive skill policy, and a margin of $mc_{M}=0.7$ when task execution is completely automated. Then, the human-machine margin $mc_{HM}$ is deterministically computed as $mc_{HM}=1-(1-mc_{H})-(1-mc_{M})=0.2$. Finally, the organization is assumed to have prior knowledge that the cost of error $c_{err}=0.9$ is incurred when execution performance falls below a threshold of $t_{err}=0.3$. The simulations are operationally configured with a single agent ($N=1$) moving through $E=10$ epochs.
Within this context, the primary objective of our decision-making agent is to evaluate the performance and net economic utility of human-exclusive ($H$), human-machine ($HM$), and machine-exclusive ($M$) skill policies. The agent also seeks to assess the impact of three alternative human-machine interaction policies: meanmean, collaborationcollaboration, and superpowersuperpower, with an applicable augmenting factor of $\gamma_{HM}=1.5$ for collaborationcollaboration and superpowersuperpower. Additionally, the study investigates the effect of two different learning curve progressions, represented by $(\delta_{H},\delta_{HM},\delta_{M})\in\{(0,0.3,0),(0,0.9,0)\}$, reflecting varying organizational abilities to improve contribution margins over time.
The aggregated simulation results, presented in Tables 1 and 2, correspond to simulation indices $s=1,\ldots,6$. The results of simulation 7 are discussed separately, but included in these tables for ease of reference.
Table 1: Simulation results: aggregated (a market perspective)
$P_{d}$ $\delta_{HM}$ $a$ $s$ $\mu(\Theta^{s})$ $\mu(Y^{s})$ $\mu(V^{s})$ $\mu(ERR^{s})$ $\mu(U^{s})$
$(1/3,1/3,1/3)$ 0.3 mean 1 0.592 0.635 0.304 0.308 -0.005
collaboratecollaborate 2 0.669 0.707 0.320 0.245 0.075
superpowersuperpower 3 0.696 0.729 0.323 0.226 0.097
0.9 mean 4 0.593 0.637 0.318 0.306 0.012
collaboratecollaborate 5 0.669 0.707 0.337 0.245 0.092
superpowersuperpower 6 0.696 0.730 0.342 0.225 0.117
$(0.1,0.4,0.5)$ 0.9 superpowersuperpower 7 0.635 0.654 0.298 0.287 0.011
The results presented in Table 1, which theoretically represent the flow of a single organization through the simulation process repeated $K$ times, can also be interpreted as the funnel of economic impact produced by a theoretical population of $K$ organizations whose skill policy decisions are uniformly distributed (same identical technological solution), detailed by the function $a$ in place. This theoretical simplification assumes that all the organizations have the same cost and margin structure, with equal access to a fair market of labor and capital, and that the technology under consideration is within reach for all of them under that same economical effort.
Based to this preamble, the aggregated results presented in Table 1 can be interpreted as a market average performance. For instance, the -0.005 economic utility observed in simulation 1 suggests a near-zero overall net benefit for the market in performing this specific task. Naturally, this is an outcome of the model that reflects its internal logics and initial setup, not an empirical survey of market performance. For this reason, this outcome is interpreted by our decisions-maker as mere theoretical reference to be compared to the results achieved by firms adopting a specific skill policy. Specifically, as we observe in Table 2, the arithmetic mean of utilities utility achieved by the firms adopting the skill policy $M$ with $\delta_{HM}=0.3$ is 0.167, remarkably over performing the market average of -0.005. In contrast, firms opting for skill policy $H$ are heavily penalized in terms of utility achieved, mainly because of the higher costs associated to a slightly lower execution performance.
Table 2: Simulation results: detail by skill policy (an organizational perspective)
$P_{d}$ $\delta_{HM}$ Skill policy ($c$) $a$ $s$ $\mu(\Theta_{c}^{s})$ $\mu(Y_{c}^{s})$ $\mu(V_{c}^{s})$ $\mu(ERR_{c}^{s})$ $\mu(U_{c}^{s})$
$(1/3,1/3,1/3)$ 0.3 H 1 0.573 0.613 0.306 0.333 -0.027
2 0.573 0.613 0.307 0.333 -0.026
3 0.573 0.612 0.306 0.334 -0.028
HM mean 1 0.592 0.645 0.149 0.306 -0.158
collaboratecollaborate 2 0.821 0.861 0.198 0.117 0.081
superpowersuperpower 3 0.902 0.927 0.213 0.060 0.154
$M\star$ 1 0.610 0.647 0.453 0.286 0.167
2 0.610 0.647 0.453 0.286 0.167
3 0.610 0.646 0.452 0.287 0.166
0.9 H 4 0.575 0.615 0.307 0.331 -0.023
5 0.571 0.609 0.305 0.336 -0.031
6 0.572 0.610 0.305 0.335 -0.029
HM mean 4 0.592 0.645 0.189 0.306 -0.117
collaboratecollaborate 5 0.821 0.860 0.251 0.118 0.132
superpowersuperpower $\star$ 6 0.899 0.925 0.269 0.061 0.208
M 4 0.613 0.651 0.456 0.283 0.173
5 0.614 0.651 0.456 0.282 0.174
6 0.613 0.651 0.455 0.282 0.173
$(0.1,0.4,0.5)$ 0.9 H 7 0.529 0.548 0.274 0.384 -0.110
HM superpowersuperpower 7 0.856 0.893 0.260 0.090 0.171
M 7 0.516 0.518 0.363 0.390 -0.027
More in general, a comparison across the first six simulation scenarios reveals a consistent improvement in net utility with the adoption of the collaboratecollaborate and superpowersuperpower deployments of human-machine interaction. While the direction of this change is predictable given the known simulation mechanics, the magnitude of the quantitative effect is less obvious. Specifically, there’s a remarkable gain when shifting from the meanmean to the collaboratecollaborate policy for human-machine skills. Furthermore, the transition from the collaborationcollaboration to the superpowersuperpower approach also yields a clear and substantial increase in performance.
Then, a more in-depth analysis of the data presented in Table 2 reveals additional local insights. When the human-machine interaction policy has a limited ability to improve its contribution margin over time ($\delta_{HM}=0.3$), the machine-exclusive skill policy offers the highest economic utility. However, the human-machine policy deployed in superpowersuperpower mode achieves a mean utility $\mu(U_{c})$ that is slightly lower. In contrast, if the $HM$ policy can significantly improve its margin of contribution ($\delta_{HM}=0.9$), the superpowersuperpower mode becomes the most economically attractive option, outperforming the machine-exclusive policy by approximately 20% in terms of mean utility. This highlights the critical role that organizational learning and improvement may have in realizing the full economic potential of human-machine collaboration.
Then, further analysis is conducted in simulation 7, where the probability distribution for task difficulty is adjusted to $P_{d}=(0.1,0.4,0.5)$, representing a higher prevalence of difficult tasks. The result, detailed in Table 2, $s=7$, show that the $HM$ skill policy achieves a mean utility of $\mu(U_{c})=0.171$. This represents the second-highest impact recorded in the entire simulation context. Notably, both human-exclusive and machine-exclusive policies fail to achieve a positive utility under these conditions.
Additionally, we may find useful to observe the evolution of economic measures through epochs. In Figure 6 we can see a compact representation of the output of simulation 3, with distributions of our metrics by skill policy (combined with the value of the function $a$), task difficulty and epoch. This allows to have a more intuitive understanding of the evolution of the economic measures of impact over time. By instance, observing the evolution of $\theta$ in this very specific case, we perceive more clearly that the distribution of values for high-difficulty tasks tends to have a slightly wider IQR for skill policy $HM$ with respect to $H$ and $M$ (intuitively, this is a direct consequence of the $\max()$ function deployed by construction by the superpowersuperpower option for the function $a$). This anecdotal finding, while irrelevant in a wider policy analysis, may be of practical interest in the specific context. More in detail, this suggests to the decision-maker that the median performance may improve when adopting the $HM$ skill policy, but at a cost of a greater variability. As a consequence, because an eventual underestimation of the cost of error $c_{err}$ in conjunction with an overestimation of $t_{err}$ could radically disrupt utility, an immediate recommendation for the agent in charge before making the skill policy decision may be to double-check these estimations with field experts.
Figure 6: Box plots of the distributions of values $\lambda_{i,c,d,e,k}$ by skill policy, task difficulty and epoch, with detail by $a$. Each row is dedicated to a metric $\lambda\in\{\theta,y,v,err,u\}$, while each column is dedicated to a specific value for the pair ($c,a$). Plots present grouping by epoch $e$ and task difficulty $d$.
This minimal example illustrates a specific scenario where a well-designed human-machine augmentation strategy can provide a competitive advantage. Then, should the task under analysis be likely to have a medium to high level of generalization difficulty in the future, the opportunity for our decision-maker is to design an effective augmentation policies to outperform competitors that either resist innovation or, on the opposite side of the spectrum, fully automate their processes. This underscores the paramount importance of correctly estimating the probability distribution of task difficulties when simulating the economic effects in production tasks and the consequential skill policy decisions.
We do not wish to enter in more detailed considerations about this example, because its scope is just to show a reasoning framework, and not to infer general conclusions through a random walk in data that is synthetic and specific. For this more general purpose, even if bounded by explicit limitations, we propose to enlarge the search space and design an experiment that may allow to approach the problem through different perspectives.
6.2 Exploring a larger space: an experimental approach
While a targeted approach (as shown in Section 6.1) is useful to address specific and known situations, a broader and more experimental approach is necessary to understand the general implications of skill policies and to navigate a larger solution space. This section details an experimental design that leverages a larger number of Monte Carlo simulations to generate comprehensive synthetic data. By moving beyond the analysis of a few individual scenarios, this methodology aims to uncover more general stylized facts about human-machine collaboration, which can then inform policy recommendations or serve as a basis for subsequent empirical validation. This approach contrasts with the prior section by prioritizing a wide-ranging exploration over a focused prediction.
6.2.1 Design of experiment
To provide the reader with a quantified measure of the effort required to explore the search space of interest, we first identify a set of discrete values for each of the variable involved (see Table 3). For categorical parameters, we limit the options to the ones discussed in Section 5, while for the numerical ones we simply split the range in a set of arbitrary steps.
Table 3: Preliminary sketch of experimental design
Variable Type Range Candidate values for simulation
$a$ Categorical $\{min,max,mean,collaborate,superpower\}$
$\gamma_{HM}$ Numeric $1,\infty]$ $\{1.1,1.3,1.5,1.7,1.9\}$
$g^{\prime}$ Categorical $\{linear,exponential,logistic,inverse\_logistic\}$
$mc_{H},mc_{HM},mc_{M}$ Numeric $[0,1]$ $\{0.1,0.3,0.5,0.7,0.9\}$
$\delta_{H},\delta_{HM},\delta_{M}$ Numeric $1,\infty]$ $\{0.1,0.3,0.5,0.7,0.9\}$
$t_{err},c_{err}$ Numeric $[0,1]$ $\{0.1,0.3,0.5,0.7,0.9\}$
Despite of the rather limited number of values for each variable, the size of a greedy exploration of all their possible combinations accounts for 39,062,500 distinct simulations. Considering that our research computational environment takes roughly 25 minutes to execute a single simulation with $N=100$, $E=10$ and $K=10000$ (no plots, only tabular data), completing the entire simulation session would take approximately 1,858 years.
Naturally, this effort is unattainable, and squeezing it in time by increasing the computational power would inevitably create a powerful contrasting force to the economic performance of the decision in the making. As a consequence, considering the typical resources available to small/medium-sized business, a compromise between effectiveness and efficiency is required to leverage the potential of this model towards informing skill policy decisions in the short run. Then, symbolically, we set our maximum allowed computational time to no more than 4 hours. To reduce the effort accordingly, we introduce three distinct simplifications.
First, we reduce the computational size of the individual simulation by setting $N=1$ and $E=10$, focusing on an individual firm walking through a minimal number of epochs, and $K=1,000$ (note: a number of replications between 5,000 and 10,000 is recommended [Mundform et al., 2011] for a second step, to consolidate the robustness of results), taking the computation time of a single simulation down to less than 5 seconds, creating room for approximately 3,000 distinct simulations.
Then, we reduce the combinatorial explosiveness of the experiment by taking action on both the categorical and numeric variables. For categorical variables, we limit the option for the function $a$ to mean, collaboratecollaborate and superpowersuperpower, and for the function $g^{\prime}$ to logistic and inverse_logistic. For the numeric variables, we introduce the following mild constraints with respect to Table 3:
    $\gamma_{HM}\in[1,2]$, thus limiting the augmenting factor to 100% increase;
    $mc_{H}\in[0.4,0.6]$ and $mc_{M}\in[0.7,0.9]$, thus putting us in the case of automation through capital allocations bringing higher margin of contribution than labor;
    $mc_{HM}=1-(1-mc_{H})-(1-mc_{M})$, to represent the idea that both labor and capital allocations are required when deploying human-machine collaboration;
    $\delta_{H}=0$ and $\delta_{M}=0$ as static default, representing the case of human and machine skills being already on their plateau of efficiency.
Then, we formulate our experimental design for numeric variables by combining the value selection optimized through the Latin Hypercube Sampling method (LHS) [McKay et al., 2000] and the Maximin method [Johnson et al., 1990] (see Figure 7 for a visualization of the search space coverage by each of the two methods).
(a) LHS method
(b) Maximin method
Figure 7: Pair plot of numeric variables selected for experimental design instructed by optimization methods
With regards to the distributions of probability: for skill policies and task difficulties we move forward with $P_{c}\sim U\{P_{H},P_{HM},P_{M}\}$ and $P_{d}\sim U\{P_{Low},P_{Med},P_{High}\}$, i.e. uniformly distributed, while for performance we opt for $P(\theta)_{c,d,e}\sim Beta(\alpha_{c,d,e},\beta_{c,d,e})$, with the parameter pairs $(\alpha,\beta)_{c,d,e=0}$ and $(\alpha,\beta)_{c,d,e=E-1}$ configuring improvement for human skills across all task difficulties, with static machine skills, as illustrated in Figure 8.
Figure 8: Distributions $P(\theta)_{c,d,e}$ for all skill policies and difficulty levels at epoch $e=0$ and e=E1e=E-1, as implemented in our first simulation batch. As we see, this setup is intended to represent human skills to evolve through epochs, while machine skills are static under the assumption of epochs to represent skill use with no re-training.
6.2.2 Results
With its relatively modest size, our experiments generates 28.8 million individual task executions, logging 144 million single data points $\theta$, $y$, $v$, errerr and $u$, individually characterized by indices $i,c,d,e,k$. To make sense of this corpus of measures, we proceed in steps.
First, we synthetically characterize the central tendency of the distributions $\omega(\Lambda^{s})$, $\omega(\Lambda_{c}^{s})$ and $\omega(\Lambda_{c,d}^{s})$ in a compact layout (see Table 4). Then, for a richer understanding of dispersion of the data, we leverage visualization based on overlapping violin and box plots. More specifically, in Figure 9 and 10 we represent the structure of the distributions $\omega(\Lambda^{s})$, $\omega(\Lambda_{c}^{s})$ and $\omega(\Lambda_{c,d}^{s})$, $\forall\Lambda$ with $\omega=\mu$ and $\omega=\sigma$, respectively. Additionally, the representations of $\omega\in\{\rho,IQR,SK\}\forall\Lambda$ are provided in Appendix, Section A.2.
Table 4: Results, $\overline{\omega(\Lambda_{(c,d)}^{s})}$ values, i.e. the arithmetic means of the distributions $\omega(\Lambda^{s})$, $\omega(\Lambda_{c)}^{s})$ and $\omega(\Lambda_{c,d}^{s})$, with $\omega\in\{\mu,\sigma,\rho,IQR,SK\}$ and $\Lambda\in\{\Theta,Y,V,ERR,U\}$
$\Lambda^{s}$ $\Lambda_{c}^{s}$ $\Lambda_{c,d}^{s}$
$\omega$ $\Theta^{s}$ $Y^{s}$ $V^{s}$ $ERR^{s}$ $U^{s}$ c $\Theta_{c}^{s}$ $Y_{c}^{s}$ $V_{c}^{s}$ $ERR_{c}^{s}$ $U_{c}^{s}$ d $\Theta_{c,d}^{s}$ $Y_{c,d}^{s}$ $V_{c,d}^{s}$ $ERR_{c,d}^{s}$ $U_{c,d}^{s}$
$\mu$ 0.65 0.64 0.35 0.09 0.26 M 0.61 0.60 0.48 0.11 0.37 High 0.33 0.33 0.27 0.24 0.03
Med 0.67 0.67 0.53 0.07 0.46
Low 0.83 0.81 0.65 0.02 0.63
HM 0.76 0.75 0.28 0.06 0.23 High 0.58 0.57 0.22 0.12 0.10
Med 0.79 0.79 0.30 0.04 0.26
Low 0.90 0.89 0.33 0.01 0.32
H 0.57 0.57 0.29 0.11 0.17 High 0.47 0.47 0.24 0.16 0.08
Med 0.56 0.56 0.28 0.11 0.17
Low 0.69 0.68 0.34 0.07 0.28
$\sigma$ 0.24 0.25 0.19 0.12 0.27 M 0.26 0.25 0.20 0.14 0.33 High 0.18 0.18 0.14 0.13 0.26
Med 0.18 0.18 0.14 0.08 0.22
Low 0.10 0.08 0.07 0.03 0.09
HM 0.19 0.20 0.09 0.08 0.16 High 0.19 0.18 0.08 0.09 0.16
Med 0.11 0.12 0.07 0.04 0.10
Low 0.06 0.06 0.05 0.02 0.07
H 0.20 0.21 0.10 0.11 0.21 High 0.20 0.21 0.11 0.13 0.23
Med 0.16 0.18 0.09 0.09 0.17
Low 0.18 0.18 0.09 0.09 0.16
$\rho$ 0.99 0.98 0.79 0.49 1.28 M 0.99 0.98 0.78 0.49 1.27 High 0.39 0.83 0.66 0.49 1.15
Med 0.88 0.83 0.66 0.40 1.07
Low 0.61 0.64 0.51 0.19 0.70
HM 0.91 0.84 0.39 0.42 0.81 High 0.85 0.79 0.37 0.41 0.78
Med 0.66 0.65 0.33 0.24 0.57
Low 0.43 0.42 0.26 0.11 0.57
H 0.97 0.93 0.47 0.48 0.95 High 0.96 0.86 0.43 0.48 0.91
Med 0.85 0.75 0.38 0.40 0.78
Low 0.90 0.84 0.42 0.42 0.84
IQRIQR 0.39 0.38 0.26 0.13 0.36 M 0.44 0.45 0.36 0.17 0.53 High 0.26 0.23 0.19 0.18 0.36
Med 0.26 0.23 0.19 0.08 0.27
Low 0.14 0.08 0.06 0.02 0.09
HM 0.24 0.24 0.12 0.06 0.19 High 0.26 0.27 0.12 0.10 0.21
Med 0.13 0.13 0.10 0.03 0.13
Low 0.07 0.05 0.08 0.01 0.10
H 0.30 0.35 0.18 0.13 0.30 High 0.31 0.38 0.19 0.16 0.36
Med 0.23 0.29 0.15 0.10 0.25
Low 0.25 0.21 0.10 0.08 0.18
SKSK -0.43 -0.23 0.49 3.93 -0.22 M -0.48 -0.40 -0.40 2.55 -0.84 High 0.47 0.41 0.41 0.54 0.25
Med -0.46 -0.41 -0.41 3.72 -1.03
Low -0.91 -1.64 -1.64 4.30 -1.96
HM -1.10 -1.16 -0.53 5.65 -1.51 High -0.18 -0.05 0.18 2.77 -0.66
Med -1.41 -2.20 -0.44 4.07 -1.34
Low -5.83 -7.00 -0.47 4.09 -1.04
H -0.23 -0.18 -0.18 3.42 -0.96 High 0.06 0.02 0.02 1.73 -0.48
Med -0.10 -0.11 -0.11 2.88 -0.82
Low -0.59 -0.54 -0.54 4.17 -1.46
Table 5: Descriptive statistics for distributions $\omega(\Lambda_{c,d}^{s})$ with detail by skill policy $c$, human-machine interaction mode $a$, and task difficulty $d$. The consultation of this descriptive representation is not straightforward and requires the effort of understanding how the meta-description works. By means of a clarifying example: the value at the intersection of the row corresponding to $\omega=IQR$, $c=HM$, $a=superpower$, $d=Med$ and the column corresponding to $\omega(U_{c,d}^{s})$ and statistic std is 0.07, which signifies that the standard deviation (std) of the values in the dataset collecting the interquartile ranges determined in each simulation $s$ for the skill policy $HM$ deployed with $a=superpower$ in the execution tasks with difficulty $d=Med$ is 0.07. A more extensive view comprising also quality-adjusted output, value and cost of error is provided in Annex, Section A.2, Table 18.
Figure 9: Main experiment, visualization of the distributions $\mu(\Lambda^{s})$, $\mu(\Lambda_{c}^{s})$ and $\mu(\Lambda_{c,d}^{s})$, with $\Lambda\in\{\Theta,Y,V,ERR,U\}$.
Figure 10: Main experiment, visualization of the distributions $\sigma(\Lambda^{s})$, $\sigma(\Lambda_{c}^{s})$ and $\sigma(\Lambda_{c,d}^{s})$, with $\Lambda\in\{\Theta,Y,V,ERR,U\}$.
7 Discussion
This section is entirely focused on the discussion of the experimental results presented in Section 6.2.2. The purpose is to move beyond the simulation of isolated scenarios towards more general insights emerging from the broader picture provided by large-scale data generated through our Monte Carlo experiment.
This section follows a two-stage road map. First, we explore a descriptive analysis of the distribution data, to identify stylized facts and emerging phenomena. Then, we pivot to meta-modeling, leveraging statistical learning methods for estimating variable importance on a computational basis. This approach allows us to move towards more general findings.
7.1 Data analysis
In first instance, our analysis assumes a uniform probability distribution for the variety of experimental configurations. This simplification facilitates a more straightforward and intuitive approach, while a subsequent weighting can be implemented in later stages of refinement to account for non-uniform probabilities.
To understand the impact of skill policy $HM$ overall, and in conjunction with the augmenting method of preference, we break down the sets $\Lambda^{s}$, $\Lambda_{c}^{s}$ and $\Lambda_{c,d}^{s}$ by the values of the function $a$. Then, for convenience and notational clarity, we define an index called HMG, denoted as $\text{HMG}_{c_{0},d,a_{0}}^{\omega,\Lambda}$, whose role is to characterize the relative gain (or loss, if negative) achieved by the skill policy $HM$ in conjunction with the function $a=a_{0}\in\{mean,collaborate,superpower\}$ with respect to skill policy $c_{0}\in\{H,M\}$, based on the mean values of the distributions $\omega(\Lambda^{s})$, $\omega(\Lambda_{c}^{s})$ and $\omega(\Lambda_{c,d}^{s})$. More formally:
[…]
Naturally, each individual HMG index provides synthetic perspectives over the central tendency of every specific distribution $\omega(\Lambda_{c,d}^{s})$. However, as the characterizing properties of the descriptors $\omega$ allow to inspect also the dispersion of the values across all the simulations in the experiment, the analysis of HMG across all the options for $\omega$ can reveal phenomena that go beyond a systematic and global averaging of all the factors taking part in the phenomena. Then, the remaining portion of this section is focused on the analysis of HMG values under the widest possible view angle.
Table 6: Analysis of $\text{HMG}_{c_{0},d,a_{0}}^{\mu,\Theta}$
Figure 11: Visualization of performance (central tendency) detailed by task difficulty and skill policy, with breakdown by augmenting effect
In a first instance, the values of $\text{HMG}_{c_{0},d,a_{0}}^{\mu,\Theta}$ presented in Table 6 and the values of $\overline{\mu(\Theta_{(c,d)}^{s})}$ represented in Figure 11 highlight that:
  • ($\star$)
    When task difficulty is High, skill policy $HM$ consistently outperforms skill policy $M$, regardless of the choice for the function $a$;
  • ($\dagger$)
    When task difficulty is Low or Med, skill policy $HM$ achieves higher performance with respect to skill policy $H$ across all $a$;
    When augmentation is achieved ($a=collaborate$ or $a=superpower$), skill policy $HM$ outperforms both skill policy $H$ and $M$ across all task difficulties;
    Even when augmentation is not achieved ($a=mean$), with low-medium difficulty task, skill policy $HM$ reaches better performance than skill policy $H$;
    Even when augmentation is not achieved ($a=mean$), skill policy $HM$ outperforms skill policy $M$ in high-difficulty tasks.
Table 7: Analysis of $HMG_{c_{0},d,a_{0}}^{\mu,U}$
Together with the synthetic overview provided by $\text{HMG}_{c_{0},d,a_{0}}^{\omega,\Lambda}$, in Figure 12 we can observe the actual distributions of $\mu(\Theta^{s})$, $\mu(\Theta_{c}^{s})$ and $\mu(\Theta_{c,d}^{s})$; this observation suggests that the comments denoted with @ and #, while correct on the mean values, can have limited validity in proximity of tails.
Figure 12: Distributions of the values in the data sets $\mu(\Theta^{s})$, $\mu(\Theta^{s}_{c})$ and $\mu(\Theta^{s}_{c,d})$, grouped by skill policy ($c$) and task difficulty ($d$)
Figure 13: Visualization of utility (central tendency) detailed by task difficulty and skill policy, with breakdown by augmenting effect
Then, the values of $\text{HMG}_{c_{0},d,a_{0}}^{\mu,U}$ presented in Table 7 and the values of $\overline{\mu(U_{(c,d)}^{s})}$ represented in Figure 13 suggest to consider that:
    Under augmentation effects, skill policy $HM$ tends to achieve better bottom-line economic results (higher utility) than skill policy $H$ over all task difficulties;
  • (ç)
    When task difficulty is High, skill policy $HM$ tends to be the preferable skill policy overall;
    When skill policy $HM$ is implemented without augmenting effects ($a=mean$), skill policy $HM$ tends to be the least economically effective skill policy option available.
The detrimental effect observed for utility in the case of failed augmentation is a direct consequence of two concurrent factors: first, the inherent cost of deploying both human and machine skills, and second the reduced ability to avoid shocks (errors) and their cost. By further focusing the analysis on alternative scenarios of cost effectiveness for machine skills within the $HM$ policy option (see Table 8, with a breakdown of $\text{HMG}_{c_{0},d,a_{0}}^{\mu,U}$ by ranges of $\delta_{HM}$), we see that an increase in the margin of contribution for $HM$ over epochs contributes to mitigate the erosion of value, while not resolving the issue. This highlights even more the importance of the augmenting effect towards the economic effectiveness of the human-machine skill policy.
Table 8: Analysis of $HMG_{c_{0},d,a_{0},\delta_{HM}}^{\mu,U}$
Figure 14: Distributions of the values in the data sets $\mu(U^{s})$, $\mu(U^{s}_{c})$ and $\mu(U^{s}_{c,d})$, grouped by skill policy ($c$) and task difficulty ($d$)
Naturally, as mentioned for $\text{HMG}_{c_{0},d,a_{0}}^{\mu,\Theta}$, these comments refer to the central tendency of the distributions involved and should not be taken as universally representative of all the scenarios. For a more complete and nuanced understanding of the distributions $\omega(U_{c,d}^{s})$, Figure 14 and Tables 9, 10, and 11 provide additional view angles. By instance, the observation of $\text{HMG}_{c_{0},d,a_{0}}^{\sigma,\Theta}$ suggests that the skill policy $HM$ contributes to reducing the central tendency of the distribution of standard deviations of utility over the experiment (data sets $U_{c,d}^{s}$). However, this tendency presents similarly across all the values of $a$, while the mean values of standard deviation computed on the sets $U_{c,d}^{s}$ tends to grow with the increase of task difficulty. The exam of other dispersion summary statistics like range and IQR ($\omega\in\{\rho,IQR\}$ confirms the effect. This observation reveals nuances which are less intuitively understandable from the model’s properties or initial conditions.
Table 9: Analysis of $HMG_{c_{0},d,a_{0}}^{\sigma,U}$
Table 10: Analysis of $HMG_{c_{0},d,a_{0}}^{\rho,U}$
Table 11: Analysis of $HMG_{c_{0},d,a_{0}}^{IQR,U}$
As our experimental results are grounded on a setup for the distributions $P(\theta)_{c,d,e}$ representing a relatively low performance of the machine-exclusive skill $M$ in high-difficulty tasks, we stress-test our findings by running the same experiment with an alternative setup for $P(\theta)_{c={M},d,e}$ -see Figure 15. A visualization of the distributions of the data resulting from the stress test is presented in Appendix, Section A.2, Figure 21 and Figure 22.
Figure 15: Distribution $P(\theta)_{c={M},d,e}$ used for the core experiment (left), vs the distribution $P(\theta)_{c={M},d,e}$ used in the stress-test experiment (right). As the reader would promptly notice, the improvement assumed for the machine skill by the stress-test setup is remarkable.
By means of a preliminary verification step, we first check that the values $\overline{\mu(U^{s}_{c_{0}=H,d})}$ in Table 6 and Table 12 are the same, as they must be; second, we check that the values of $\overline{\mu(U^{s}_{c_{0}=M,d})}$ and $\overline{\mu(U^{s}_{c_{0}=HM,d})},\forall a$ are increased with respect to the core experiment, as they intuitively should. This result is an obvious consequence of the new setup, and deserves no particular comment. Instead, it is more interesting to look at Tables 13 and 15, where we can appreciate the impact of the stress-test on central tendency and dispersion of utility over the simulations.
Table 12: Stress test with high machine skills, Analysis of $HMG_{c_{0},d,a_{0}}^{\mu,\Theta}$
The data presented in Table 13 suggests that the skill policy $M$ now becomes the single most attractive option from the economical perspective across all task difficulties, with no augmenting effects being able to close the gap between the human-machine skill policy versus the machine-exclusive one. However, as detailed in Table 14 and illustrated in Figure 16, focusing the lens on situations characterized by potentially high cost of error, the human-machine skill policy $HM$ featuring augmenting effects still prevails in terms of utility over the machine exclusive policy $M$ when the task difficulty is high. This reinforces the idea that, whatever the performance achieved by artificial skills, the factors determining the occurrence of errors and the magnitude of their consequences (in our case, quantified by economic metrics) impose a cap to the degree of automation that can be achieved by a responsible agent.
Table 13: Stress test with high machine skills, Analysis of $HMG_{c_{0},d,a_{0}}^{\mu,U}$
Table 14: Stress test with high machine skills, Analysis of $HMG_{c_{0},d,a_{0}}^{\mu,U}$ in conditions characterized by potential high cost of errors
Figure 16: Distributions of the values in the data sets $\mu(U^{s})$, $\mu(U^{s}_{c})$ and $\mu(U^{s}_{c,d})$, grouped by skill policy ($c$) and task difficulty ($d$), in our stress test, with focus on situations characterized by high cost of errors ($t_{err}\leq 0.1$ and $c_{err}\geq 0.80$)
Finally, a review of the Table 15 confirms that the $HM$ policy tends to reduce the dispersion of utility values with respect to skill policies $H$ and $M$, with augmentation strengthening this effect.
Table 15: Stress test with high machine skills, Analysis of $HMG_{c_{0},d,a_{0}}^{\sigma,U}$
7.2 Meta-modeling
To understand more about the importance of the model’s variables on performance and utility, we leverage the potential of introspection analysis on meta-models (more specifically, surrogate models) learned on our synthetic data by means of i) Permutation Importance analysis and ii) SHapley Additive exPlanations method (SHAP).
Permutation Importance analysis [Breiman, 2001a] is a computational method measuring the importance of a feature by determining the increase of the prediction error after the values of that feature are randomly permuted (a larger increase in error denotes a more important feature). In the context of our work, we apply permutation analysis for each of the models trained in cross validation, finally computing mean and standard deviation of the results; also, we aggregate categorical values encoded in one-hot variables in a unique feature, improving interpretability.
In contrast, the SHAP method [Štrumbelj and Kononenko, 2014] [Datta et al., 2016] [Lundberg and Lee, 2017], named by the acronym of SHapley Additive exPlanations, computes the marginal contribution provided by each feature to the output for any given example, providing local interpretability. Then, global model explanations are obtained by aggregating Shapley values across all examples in the data set, suggesting which features are most influential overall, and how they impact the model’s predictions.
As the features $c$ and $a$ have a strong correlation by construction ($a$ assumes value individualindividual when $c\neq HM$ and either meanmean, collaboratecollaborate or superpowersuperpower when $c=HM$), to avoid compromising the response of permutation analysis and SHAP [Strobl et al., 2008] [Lundberg and Lee, 2017], we combine these features in a unique new variable called c_ac\_a. Also, we drop the variables $delta_{H}$ and $delta_{M}$ because, by construction in this specific implementation, they are constant. Additionally, we select all the variables $mc_{c}$ because, although their individual range is constrained to fit real-world cost dynamics, they are free to combine each-other without any direct relationship, thus relaxing the condition set in the small-scale exploration illustrated in Section 6.1. The adjusted Generalized Variance Inflation Factor (Adj. GVIF) [Fox and Monette, 1992] for the new set of variables (see Table 17, as the result of GVIF is model-dependent) confirms the absence of correlation, thus allowing to proceed firmly.
With regards to the modeling strategy, we implement Random Forest regressors (RFr) and Gradient Boosting regressors (GBr), measuring their performance through $MSE$ and $R^{2}$ (the purpose of measuring the predictive performance is to make sure that the model is effectively compressing data, and learning the information encoded in it, making the interpretation of variable importance more reliable). To give the necessary robustness to the propaedeutic performance assessment, we execute a 10-Fold Cross Validation (CV), removing duplicated examples from the dataset which is designated to the meta-model of interest.
According to the provisions illustrated, the following four models are learned, and their respective performance presented in Table 16:
  1. 1.
    $\hat{f}:(c\_a,d,\gamma_{HM})\rightarrow\mu(\Theta_{c,d})$, denoted as MM1; this model is aimed at predicting $\mu(\Theta_{c,d})$ given the skill/augmentation policy c_ac\_a, the task difficulty $d$, and the augmenting factor $\gamma_{HM}$;
  2. 2.
    $\hat{f}:(c\_a,\gamma_{HM})\rightarrow\mu(\Theta_{c})$, denoted as MM2, through which the skill level by category is predicted only given c_ac\_a, and $\gamma_{HM}$;
  3. 3.
    $\hat{f}:(c\_a,d,\gamma_{HM},g^{\prime},mc_{c},\delta_{HM},t_{err},c_{err})\rightarrow\mu(U_{c,d})$, denoted as MM3, aimed at predicting $\mu(U_{c,d})$ given c_ac\_a, $d$, $\gamma_{HM}$, $g^{\prime}:\theta\rightarrow y$, the tuple $mc_{c}$, $\delta_{HM}$, and the pair ($t_{err},c_{err}$);
  4. 4.
    $\hat{f}:(c\_a,\gamma_{HM},g^{\prime},mc_{c},\delta_{HM},t_{err},c_{err})\rightarrow\mu(U_{c})$, denoted as MM4.
Table 16: Results of the 10-Fold Cross Validation (MSE, $R^{2}$)
RFr GBr
Meta-model MSE $R^{2}$ MSE $R^{2}$
MM1: $\mu(\Theta_{c,d})=\hat{f}(c\_a,d,\gamma_{HM})$ 0.0000 0.9994 0.0006 0.9842
MM2: $\mu(\Theta_{c})=\hat{f}(c\_a,\gamma_{HM})$ 0.0000 0.9989 0.0003 0.9731
MM3: $\mu(U_{c,d})=\hat{f}(c\_a,d,\gamma_{HM},g^{\prime},mc_{c},\delta_{HM},t_{err},c_{err})$ 0.0003 0.9940 0.0039 0.9319
MM4: $\mu(U_{c})=\hat{f}(c\_a,\gamma_{HM},g^{\prime},mc_{c},\delta_{HM},t_{err},c_{err})$ 0.0002 0.9910 0.0019 0.9247
Since the models’ performance reveals to be capturing the signal in the data, we move forward with introspection analyses. The results of Permutation Importance, summarized in Table 17, and the SHAP swarm plots represented in Figures 17, 18, 19 and 20 show a remarkably consistent response of the two methods, allowing consideration.
Table 17: Results of Permutation Importance analysis (10 cross validation iterations) and adjusted GVIF.
Mean Decrease of $R^{2}$
Meta-model Target Feature Adj. GVIF RFr GBr
mean std mean std
MM1 $\mu(\Theta_{c,d})$ $d$ 1.00 1.280 0.044 1.270 0.047
c_ac\_a 1.00 0.859 0.036 0.806 0.016
$\gamma_{HM}$ 1.00 0.094 0.005 0.060 0.005
MM2 $\mu(\Theta_{c})$ c_ac\_a 1.00 1.954 0.086 1.871 0.060
$\gamma_{HM}$ 1.00 0.232 0.024 0.165 0.020
MM3 $\mu(U_{c,d})$ $d$ 1.00 1.456 0.065 1.039 0.035
c_ac\_a 1.00 0.784 0.027 0.612 0.015
$t_{err}$ 1.00 0.251 0.013 0.209 0.009
$c_{err}$ 1.01 0.223 0.009 0.175 0.010
$mc_{HM}$ 1.00 0.142 0.007 0.085 0.007
$g^{\prime}$ 1.00 0.096 0.004 0.067 0.003
$\gamma_{HM}$ 1.00 0.023 0.003 0.006 0.001
$mc_{M}$ 1.01 0.015 0.001 0.008 0.001
$\delta_{HM}$ 1.00 0.012 0.001 0.006 0.001
$mc_{H}$ 1.00 0.010 0.000 0.007 0.001
MM4 $\mu(U_{c})$ c_ac\_a 1.00 1.109 0.063 0.933 0.070
$t_{err}$ 1.00 0.480 0.025 0.426 0.035
$c_{err}$ 1.01 0.369 0.041 0.322 0.027
$mc_{HM}$ 1.00 0.308 0.023 0.210 0.016
$g^{\prime}$ 1.00 0.048 0.002 0.048 0.006
$\gamma_{HM}$ 1.00 0.040 0.004 0.014 0.002
$mc_{M}$ 1.01 0.027 0.002 0.018 0.003
$mc_{H}$ 1.00 0.023 0.002 0.018 0.001
$\delta_{HM}$ 1.00 0.021 0.002 0.014 0.003
For the meta-model 1 (MM1), the analysis illustrated in Figure 17 suggests that, towards predicting the task-execution performance by skill policy and task difficulty, the highest importance is deserved to task difficulty $d$, followed by the skill/augmentation policy c_ac\_a, with relatively marginal contribution of $\gamma_{HM}$.
The response of SHAP for MM2 (see Figure 18) confirms the importance of the skill/augmentation policy c_ac\_a in terms of impact on $\mu(\Theta_{c})$, which is also highlighted in the analysis conducted on meta-models MM3 and MM4. Moreover, the application of SHAP to MM3 confirms the strong impact of task difficulty also on utility. Finally, jointly observing the results of SHAP on MM3 and MM4 denotes the impact of skill/augmentation policy on utility, as well as the contribution of the variables parameterizing the logics of cost of errors ($t_{err}$ and $c_{err}$).
(a) Random Forest regressor
(b) Gradient Boosting regressor
Figure 17: MM1, SHAP summary (beeswarm) plot illustrating feature impact on the model output. Features are ranked by global importance. Each point represents the SHAP value for a specific instance, with the color indicating the feature’s value. Categorical values (extremes): for task difficulty $d$, blue dots correspond to Low and red dots to High; for skill/augmentation policy c_ac\_a, blue dots correspond to H_individual and red dots to HM_superpower, with intermediate color scale passing through HM_mean, M_individual, and HM_collaborate.
(a) Random Forest regressor
(b) Gradient Boosting regressor
Figure 18: MM2, SHAP summary (beeswarm).
(a) Random Forest regressor
(b) Gradient Boosting regressor
Figure 19: MM3, SHAP summary (beeswarm). Categorical values (extremes): for function $g^{\prime}$, the blue dots represent logistic, while red dots represent inverse_logistic.
(a) Random Forest regressor
(b) Gradient Boosting regressor
Figure 20: MM4, SHAP summary (beeswarm) plot.
The combined observation of the results given by Permutation Importance and SHAP are consistent with our exploratory data analysis, quantitative estimating the importance of the skill/augmentation policy and of the dynamics related to the cost of errors towards both performance and utility (task difficulty is essential to break down the impact).
Interestingly, the baseline margin of contribution associated to skill policy $HM$ results remarkably more important that the corresponding value related to skill policy $H$ or $M$. In contrast, the pace of growth of the margin ($\delta_{HM}$) reveals to be a lower-importance feature. Equally remarkable is the relatively low importance of $g^{\prime}$ and $\gamma_{HM}$ (the reduce importance of the augmentation gain versus the augmentation skill/augmentation policy is interesting, suggesting that the multiplicative factor itself plays a minor role with respect to the skill interaction function).
7.3 Limitations
Combining Monte Carlo simulations with machine learning meta-models can provide valuable insights, but limitations must be acknowledged to contextualize findings appropriately.
The primary limitation of this study is that its conclusions are derived from synthetically generated data. Essentially, the simulation creates an idealized in silico environment whose potential gaps with the real-world should always be considered with the highest regard, as complex or unmodeled interactions could fundamentally alter the outcomes [Law et al., 2007]. As a consequence, the validity of our findings is contingent upon the fidelity of our model, which is governed by two sets of assumptions:
    Design hypotheses, related to the theoretical mechanics and interactions encoded in the model;
    Setup hypotheses, concerning the specific characterization of the scenarios of interest through parameters and probability distributions chosen for the stochastic elements; then, before verification and validation, results and insights should be intended as specific to the explored parameter space.
Design hypotheses include the specific choice in terms of distributions (for example, the choice of $P(\theta)_{c,d,e}\sim Beta(\alpha_{c,d,e},\beta_{c,d,e})$) and the logics of computation of economic measures (for example, the cost of errors). Another arbitrary design choice pertains the specific integration of stochastic effects, which in our work are limited to the realization of task-execution performance, leaving the other measures of economic impact to deterministic computation. This can potentially inject a re-scaling effect of the performance distribution down through the funnel of economic value, thus making the dynamics of bottom-line measures (for example, utility) more a direct reflection of performance rather than an emergent behavior (this may be especially considered when evaluating tail risks).
With regards to setup hypotheses, the range of values considered for our first experiment may be limiting, especially for what concerns alternative choices for the parameters determining the Beta distributions, or the choice of other classes of distributions altogether. Also, the number of replications we deployed for each simulation ($K=1,000$) is relatively low, suggesting the importance to plan further experiments to validate the current results and the related findings. As a consequence, our findings should be considered bounded to the setup conditions, with special regard to the parameters that are constant over the experiment. However, the approach offers the opportunity to resolve this specific issue by scaling resources in width and depth towards a more extensive and robust exploration of the search space.
A more extensive treatment of the potential limitations of the simulation approach, and the corresponding mitigating actions, is provided in Section 5.2.
Moreover, we underline that Permutation Importance and SHAP are not meant to reveal causal relationship between features and outcome, but importance with respect to the predictions computed by the model under consideration.
Additionally, we define economic utility purely in terms of net monetary gain. While this approach reveals the focus on measures of immediate understanding and utility for the organizational agent, it may provide an incomplete picture from a the perspective of social sciences.
8 Conclusions
With this work, we intend to offer a contribution to the economic-minded decision-makers confronting with the problem of skill policy decisions in the context of cognitive tasks characterized by a spectrum of generalization difficulty.
More specifically, we address two challenges: first, the estimation of the quantitative economic impact of the options of choice when navigating relatively small search spaces, leveraging a limited number of situations and simulations to produce distributions of synthetic data that can be quickly analyzed; second, in the context of larger search spaces representing higher levels of uncertainty, the design of experiments involving a large number of simulations whose result can allow to find more general insights and policy implications.
In this frame of reference, we model task execution performance as a non-deterministic outcome of complex interactions between skills and uncertain task content, crystallizing the idea of measuring the economic impact of skill policy decisions from the bottom-up. By combining the provisions of the seminal framework illustrated in [Farboodi and Veldkamp, 2021] with the concepts presented in [Chollet, 2019], we define a theoretical model and a practical Monte Carlo simulation environment to generate synthetic representations of the economic impact of alternative skill policies involving human and/or machine skills confronting with varying levels of generalization difficulty.
Our experimental setup establishes a baseline characterizing a specific stage of technological maturity, using Beta distributions to model machine skills that excel at tasks of low-to-medium difficulty while being less effective on complex tasks demanding high levels of generalization. This performance profile for machine skills, which are set to be static once trained, is combined with human skills that evolve over epochs, partially closing the performance gap with machine skills on low and medium-difficulty tasks, and widening their advantage on high-difficulty tasks. The margin of contribution for automation is set to be empirically higher than for human labor to reflect lower variable costs and higher economies of scale. Human-machine interaction, initially floating in a range compatible with the sum of the costs of both human and machine skills, improves its margin of contribution over epochs up to a factor of 2x. No constraint is set for the combinations of the minimum performance triggering an error shock and its unitary cost. This setup, combined with non-linear performance-to-output relationships, contribute to characterize an experimental scenario that is grounded in empirical pragmatism, suitable to representing a common real-world case.
In this context of reference and specific conditions, our experimental results highlight the following stylized facts:
    Regardless of task difficulty, a human-machine skill policy that facilitates augmentation tends to outperform a human-only approach, with an observed advantage in terms of utility ranging from +25% to +175% (central tendency);
    Under conditions of high task difficulty, augmenting human-machine policies emerge as economically impactful, with an observed utility advantage (central tendency) between +50% to +175% versus human-exclusive skill policy, and between four to seven times that of a machine-exclusive skill policy;
    A strategy that combines human and machine skills but fails to produce augmenting effects tends to be the least economically favorable option. This approach results in a significant erosion of utility, from -11% in low-difficulty tasks to a -233% in high-difficulty tasks.
A second experiment, aimed at stress-testing our findings in a setup characterized by machine skills achieving high performance across all task difficulties, confirms that automation can progressively extend its economic effectiveness to a wider set of situations as the machine skills improve. However, the human-machine skill policy still tends to prevail in high-difficulty task when the cost of error is high. Also, both our experiments show that the combination of human and machine skills tends to reduce the dispersion of values for utility, and that this reduction is more evident when augmenting effects are involved.
Finally, we learn Random Forest and Gradient Boosting regression meta-models on our core experimental data to further analyze the relationship between the model variables and responses. To this end, we deploy Permutation Importance tests and SHAP. The results of the introspective analysis provide additional support to the idea that task difficulty, skill policy and the level of augmentation achieved are the most important predictors for economic impact within the proposed simulation environment and setup. Also important are the the factors conditioning the emergence of economic consequences of errors, and the baseline margin of contribution associated with skill policies based on human-machine interaction.
Then, the primary takeaway for organizational leaders and policymakers is that in tasks characterize by high difficulty and high cost of error, a human-machine skill policy delivering strong augmenting effects represents an opportunity to achieve better economic results with respect to the human-exclusive or machine-exclusive options. However, without augmenting effects, the inherent costs of the human-machine model can make it less efficient, and the increasing costs due to errors can drag the net economic impact of the dual-skill policy down to zero, or even in a negative danger zone. As a consequence, a strategy based on human-machine integration is not a magic button nor a risk-mitigating solution, but a strategic entrepreneurial option of choice that requires full commitment to generate a competitive advantage.
At the same time, our core results quantitatively support the established notion that automation tends to be the most effective choice for low-to-medium difficulty tasks, with a progressive expansion of the boundaries of economical effectiveness as machine skills improve their performance in task characterized by medium-to-high difficulty. However, when tasks combine high difficulty with high cost of errors, a human-machine skill policy carefully designed for augmentation still represents the most effective option for the responsible and economic-minded agent.
A second-level takeaway for decision-makers pertains the importance of knowing as better as possible the process of interest, and the characteristics of the environment in which the task unfolds. More specifically, being able to obtain high-quality estimates of the probability distribution of difficulty for the task of interest, together with robust assessments of the performance of the candidate skills by task difficulty, can allow to leverage the potential of simulation towards better supporting skill policy decisions.
By modeling the skill policy decision problem through an in-silico simulation environment able to replicate stylized facts already established in the literature, we provide a promising and scalable framework for organizations to test alternative skill policies and their potential economic outcomes without incurring real-world costs. The ability to model and forecast the economic impact of skill policy options in advance can be, itself, a competitive advantage, enabling organizations to make more informed, efficient, and profitable strategic decisions in an era of unprecedented technological change.
Naturally, this poses an additional challenge for the modern organization: develop the ability to describe complex processes and their environment through the lens of data by means of a structured and systematic method that can gradually become part fo the organizational administrative behavior.
9 Future developments
To mitigate the limitations of the in-silico environment created by simulation, empirical validation would radically improve and strengthen this work.
Additionally, as acknowledged in Section 7.3, as the model is apt to deal with multiple stochastic effects, promising future developments may involve specif treatments aimed at modeling the unitary margin of contribution $mc_{c}$ as probability distributions to reflect market volatility or, simply, the variability of cost structures. Similarly, the cost of an error ($c_{err}$) could be modeled as a distribution to accommodate cases where the penalty suffered for a failed task cannot be easily traced back to a deterministic number. Also, exploring other setups for the parameters $\alpha_{c,d,e}$ and $\beta_{c,d,e})$ of $P(\theta)_{c,d,e}\sim Beta(\alpha_{c,d,e},\beta_{c,d,e})$ may be a promising option for future developments of this work, together with exploring options beyond the $Beta$ distribution, as well as increasing the number of repetition in the Monte Carlo process.
Moreover, the evolution of skills through linear interpolation may be refined to consider other more complex and nuanced dynamics, more in line with the true nature of learning and adaptation. A more advanced approach could model the evolution of skills as a non-linear function of cumulative experience, or data generated. This is also in line with the provisions illustrated in [Farboodi and Veldkamp, 2021], where data is a defined as a by-product of production potentially enabling and feeding a feedback loop. Also, future designs may specifically model dynamic training for machine skills, to account for the differences distinguishing the training of statistical models from heuristics or, more in general, of the models operated by the human mind.
10 Acknowledgments
We thank Prof. Enrico Minelli and Prof. Carmen Marchiori for the stimulating discussion about data economy in the early phases of this study. AI tools have been used for light copyediting on small portions of this manuscript. All errors are our own. No conflict of interest involving the author exist at the time this draft is published.
Appendix A Appendix
A.1 Algorithms
Algorithm 1 Function randomchoice()
[…]
2:$n$
[…]
Return: choicechoice $\triangleright$ A vector of length $n$
Algorithm 2 Function randomsample()
[…]
3:$n$
[…]
Return: samplesample $\triangleright$ A vector of length $n$
Algorithm 3 Function interpolate()
1:param_startparam\_start
2:param_endparam\_end
3:cur_stepcur\_step
4:tot_stepstot\_steps
5:if $total\_steps\leq 1$ then
6:   $cur\_param\leftarrow param\_start$
7:else
8:   $cur\_param\leftarrow param\_start+(param\_end-param\_start)*cur\_step/(tot\_steps-1)$
9:end if
Return: cur_paramcur\_param
Algorithm 4 Simulation
[…]
10:$K,N,E$
[…]
15:for $k=1$ to $K$ do
16:   for $e=0$ to E1E-1 do
17:    if $e=0$ then
18:      $c\leftarrow randomchoice(C,N,P_{c})$
19:    end if
20:    for $i=0$ to N1N-1 do
21:      $d\leftarrow randomchoice(D,1,P_{d})$
22:      $\alpha_{c,d,e}\leftarrow interpolate(\alpha_{c,d,e=0},\alpha_{c,d,e=E-1},e,E)$
23:      $\beta_{c,d,e}\leftarrow interpolate(\beta_{c,d,e=0},\beta_{c,d,e=E-1},e,E)$
24:      $\theta_{i,c=H,d,e,k}\leftarrow randomsample\big(\alpha_{c=H,d,e},\beta_{c=H,d,e},1)$
25:      $\theta_{i,c=M,d,e,k}\leftarrow randomsample\big(\alpha_{c=M,d,e},\beta_{c=M,d,e},1)$
26:      if $c=H$ then
27:       $\theta_{i,c,d,e,k}\leftarrow\theta_{i,c=H,d,e,k}$
28:      else if $c=M$ then
29:       $\theta_{i,c,d,e,k}\leftarrow\theta_{i,c=M,d,e,k}$
30:      else if $c=HM$ then
31:       $\theta_{i,c,d,e,k}\leftarrow a(\theta_{i,c=H,d,e,k},\theta_{i,c=M,d,e,k},\gamma_{HM})$
32:      end if
33:      $y_{i,c,d,e,k}\leftarrow g^{\prime}(\theta_{i,c,d,e,k})$
34:      $\forall c:\delta_{c,e}\leftarrow interpolate(1,1+\delta_{c},e,E)$
35:      $v_{i,c,d,e,k}\leftarrow y_{i,c,d,e,k}*mc_{c}*\delta_{c,e}$
36:      if $(1-\theta_{i,c,d,e,k})\geq t_{err}$ then
37:       $err_{i,c,d,e,k}\leftarrow(1-\theta_{i,c,d,e,k})*c_{err}$
38:      else
39:       $err_{i,c,d,e,k}\leftarrow 0$
40:      end if
41:      $u_{i,c,d,e,k}\leftarrow v_{i,c,d,e,k}-err_{i,c,d,e,k}$
42:    end for
43:   end for
44:end for
Result: A generic simulation $s$ returns the sets $\Lambda^{s}=\{\lambda_{i,c,d,e,k}^{s}\}$ with $\lambda\in\{\theta,y,v,err,u\}$:
    $\Theta^{s}=\{\theta_{i,c,d,e,k}^{s}\}$;
    $Y^{s}=\{y_{i,c,d,e,k}^{s}\}$;
    $V^{s}=\{v_{i,c,d,e,k}^{s}\}$;
    $ERR^{s}=\{err_{i,c,d,e,k}^{s}\}$;
    $U^{s}=\{u_{i,c,d,e,k}^{s}\}$.
A.2 Results, additional material
Table 18: Main experiment, visualization of the distributions $\omega(\Lambda_{c,d}^{s})$ with detail by skill policy $c$, human-machine interaction mode $a$, and task difficulty $d$.
Figure 21: Stress-test experiment, visualization of the distributions $\mu(\Lambda^{s})$, $\mu(\Lambda_{c}^{s})$ and $\mu(\Lambda_{c,d}^{s})$, with $\Lambda\in\{\Theta,Y,V,ERR,U\}$.
Figure 22: Stress-test experiment, visualization of the distributions $\sigma(\Lambda^{s})$, $\sigma(\Lambda_{c}^{s})$ and $\sigma(\Lambda_{c,d}^{s})$, with $\Lambda\in\{\Theta,Y,V,ERR,U\}$.
Figure 23: Main experiment, visualization of the distributions $\rho(\Lambda^{s})$, $\rho(\Lambda_{c}^{s})$ and $\rho(\Lambda_{c,d}^{s})$, with $\Lambda\in\{\Theta,Y,V,ERR,U\}$.
Figure 24: Main experiment, visualization of the distributions $IQR(\Theta_{(c,d)}^{s})$, $IQR(Y_{(c,d)}^{s})$, $IQR(V_{(c,d)}^{s})$, $\mu(ERR_{(c,d)}^{s})$ and $IQR(U_{(c,d)}^{s})$.
Figure 25: Main experiment, visualization of the distributions $SK(\Theta_{(c,d)}^{s})$, $SK(Y_{(c,d)}^{s})$, $SK(V_{(c,d)}^{s})$, $\mu(ERR_{(c,d)}^{s})$ and $IQR(U_{(c,d)}^{s})$.
A.3 EDA, additional material
Figure 26: Visualization of the distributions characterizing the result of an individual simulation, small search space (simulation 1/7)
Figure 27: Visualization of the distributions characterizing the result of an individual simulation, small search space (simulation 2/7)
Figure 28: Visualization of the distributions characterizing the result of an individual simulation, small search space (simulation 3/7)
Figure 29: Visualization of the distributions characterizing the result of an individual simulation, small search space (simulation 4/7)
Figure 30: Visualization of the distributions characterizing the result of an individual simulation, small search space (simulation 5/7)
Figure 31: Visualization of the distributions characterizing the result of an individual simulation, small search space (simulation 6/7)
Figure 32: Visualization of the distributions characterizing the result of an individual simulation, small search space (simulation 7/7)
Table 19: Simulation results: detail by skill policy and task difficulty
$P_{d}$ $\delta_{HM}$ $s$ Skill policy ($c$) $a$ Task difficulty ($d$) $\mu(\Theta_{c,d}^{s})$ $\mu(Y_{c,d}^{s})$ $\mu(V_{c,d}^{s})$ $\mu(ERR_{c,d}^{s})$ $\mu(U_{c,d}^{s})$
$(1/3,1/3,1/3)$ 0.3 1 H Low 0.685 0.775 0.387 0.202 0.185
Med 0.558 0.599 0.300 0.356 -0.056
High 0.475 0.464 0.232 0.442 -0.210
HM mean Low 0.760 0.898 0.207 0.091 0.116
Med 0.611 0.705 0.163 0.297 -0.135
High 0.406 0.331 0.077 0.531 -0.454
M Low 0.832 0.942 0.660 0.039 0.621
Med 0.665 0.749 0.525 0.225 0.300
High 0.334 0.251 0.175 0.594 -0.418
2 H Low 0.687 0.778 0.389 0.200 0.189
Med 0.558 0.599 0.299 0.356 -0.057
High 0.476 0.463 0.232 0.442 -0.210
HM collaboratecollaborate Low 0.981 0.990 0.227 0.002 0.226
Med 0.879 0.946 0.218 0.042 0.175
High 0.605 0.649 0.150 0.306 -0.155
M Low 0.832 0.942 0.660 0.039 0.621
Med 0.665 0.749 0.524 0.226 0.298
High 0.334 0.251 0.176 0.593 -0.417
3 H Low 0.683 0.773 0.386 0.204 0.182
Med 0.560 0.601 0.300 0.354 -0.053
High 0.475 0.462 0.231 0.443 -0.211
HM superpowersuperpower Low 0.998 0.993 0.228 0.000 0.228
Med 0.947 0.974 0.224 0.017 0.208
High 0.758 0.812 0.187 0.164 0.024
M Low 0.833 0.943 0.660 0.039 0.621
Med 0.664 0.746 0.523 0.227 0.296
High 0.332 0.248 0.174 0.595 -0.421
0.9 4 H Low 0.687 0.778 0.389 0.198 0.191
Med 0.559 0.600 0.300 0.356 -0.056
High 0.475 0.462 0.231 0.442 -0.211
HM mean Low 0.758 0.898 0.262 0.091 0.171
Med 0.613 0.709 0.207 0.295 -0.087
High 0.402 0.325 0.096 0.534 -0.438
M Low 0.833 0.943 0.660 0.037 0.623
Med 0.665 0.749 0.524 0.225 0.299
High 0.339 0.257 0.180 0.588 -0.408
5 H Low 0.685 0.776 0.388 0.202 0.185
Med 0.557 0.596 0.298 0.357 -0.059
High 0.471 0.457 0.228 0.447 -0.219
HM collaboratecollaborate Low 0.981 0.989 0.287 0.002 0.285
Med 0.882 0.946 0.275 0.042 0.232
High 0.605 0.649 0.191 0.307 -0.116
M Low 0.834 0.943 0.660 0.038 0.623
Med 0.668 0.753 0.527 0.221 0.306
High 0.336 0.253 0.177 0.591 -0.414
6 H Low 0.685 0.774 0.387 0.201 0.186
Med 0.560 0.602 0.301 0.354 -0.053
High 0.471 0.457 0.228 0.447 -0.219
HM superpowersuperpower Low 0.998 0.993 0.289 0.000 0.289
Med 0.947 0.975 0.283 0.016 0.267
High 0.756 0.810 0.237 0.165 0.072
M Low 0.835 0.944 0.661 0.038 0.623
Med 0.666 0.749 0.524 0.223 0.301
High 0.336 0.253 0.177 0.591 -0.413
$(0.1,0.4,0.5)$ 0.9 7 H Low 0.684 0.772 0.386 0.202 0.184
Med 0.557 0.596 0.298 0.357 -0.059
High 0.476 0.465 0.232 0.441 -0.209
HM superpowersuperpower Low 0.998 0.993 0.288 0.000 0.288
Med 0.947 0.974 0.283 0.017 0.266
High 0.756 0.810 0.237 0.165 0.071
M Low 0.831 0.941 0.659 0.040 0.618
Med 0.669 0.755 0.529 0.219 0.309
High 0.332 0.247 0.173 0.595 -0.423
A.4 From data to skill policies
This manuscript assumes that all organizations have equal access to the same skill programs, making no specific mention of the method they use to determine their preferential skill $H$ and $M$ for the task of interest. This simplification does not impact the generality of our work, but deserves a more detailed treatment to better connect all the pieces of the puzzle, towards practical examples.
In a more general setting, at any arbitrary time $t$ characterized by a specific state of technological advancement, given the task of interest $z_{i}$ and its description $D_{i,t}$, the decision-maker confronts with the following problem:
$\hat{Sol}_{i,t}=\operatorname*{arg\,max}_{j,c}{\bigg(E\Big[U({Sol}_{i,t,c,j})\Big]\bigg)}$ (17)
where $E\big[U({Sol}_{i,t,c,j})\big]$ is the expected utility achieved by deploying the skill ${Sol}_{i,t,c,j}$, where $j$ denotes a specific implementation (or instance) of a skill [^2] The notational characterization of a generic technological solution as $Sol$ is derived from [Chollet, 2019], in which $Sol_{T}^{\theta}$ denotes the shortest solution for the task $T$ achieving performance $\theta$ at evaluation time.. For example, in case of a heuristic ($c=H$), $j$ may represent different procedures, or different operators; in contrast, for statistical models ($c=M$), $j$ represents different and alternative classes of models (a neural network, or random forest, or a linear regression), as well as alternative implementations of the same model.
Assuming that the informational set $D_{i,t}$ is not equally accessible or usable by agents to define the solution of choice (for example, only a subset of $D_{i,t}$ may be accessible or processable by the human mind due to bounded computational capabilities), we first define $D_{i,t,c}\subseteq D_{i,t}$ the subset of information accessible to an agent developing a skill of category $c\in\{H,M\}$. Then, $D_{i,t,c}$ is split in a training curriculum $C_{i,t,c}$ and an evaluation curriculum $E_{i,t,c}$, with $C_{i,t,c}\cup E_{i,t,c}\subseteq D_{i,t,c}$ and $C_{i,t,c}\cap E_{i,t,c}=\emptyset$.
When no perturbation of $C_{i,t,c}$ and $E_{i,t,c}$ is required, the candidate solutions ${Sol}_{i,t,c,j}$ are trained on $C_{i,t,c}$ by minimizing the loss function of choice. Upon evaluation on the hold-out set $E_{i,t,c}$, each solution produces a distribution of results in terms of performance $\theta_{i,t,c,j}$, characterized by the probability distribution $P(\theta)_{i,t,c}$. When it is possible to decompose $E_{i,t,c}$ in disjoint subsets $E_{i,t,c,d}$, where $d$ is the level of generalization difficulty characterizing each example, it is possible to determining $P(\theta)_{i,t,c,d}$. Once we drop the index $i$ (in our simulation environment we focus on a single task) and we take the perspective of task evolution over epochs denoted with $e$, these distributions are exactly the $P(\theta)_{c,d,e}$ required as input by our model. Then, the expectations $E\big[U({Sol}_{i,t,c,j})\big]$ are computed by simulation, and the problem can be resolved, confronting the desired set of options for the augmenting function $a$ (in an alternative framing, we may consider $a$ as an additional variable for the argument of the maxima instead of having alternative options for $a$ to populate a distribution of results; these choices may be transparent to the process, and left to individual preference).
Figure 33: Schematic representation of the process of learning and performance assessment of a skill program $Sol_{i,t,c,j}$, where $j$ denotes an alternative candidate model instance. While a final model is learned on the entire accessible data set $D_{i,t,c}$, the performance assessment aimed at estimating the distribution $P(\theta)$ requires to perform training and evaluation (test) according to the method of choice, which may involve multiple rounds of training/evaluation based on perturbed data sets (each iteration denoted with index $q=1,\ldots,Q$. The final distribution $P(\theta)$ serves as input for the simulation, which returns the estimation of economic impact for $Sol_{i,t,c,j}$.
In Figure 33 we illustrate a schematic representation of the process, generalizing to the case of perturbations of the training/evaluation set (for example, by deploying cross-validation). As we see, the distributions $P(\theta)_{c[,e]}$ serving as input for the simulation process do not come from thin air, but they can be estimated through a rigorous process that may involve multiple candidate skill program instances. In Figure 34 we describe the additional complexity of estimating $P(\theta)_{c,d[,e]}$ when it is possible to decompose $E_{i,t,c[,q]}$ by task difficulty $d$. In this representation, we consider producing a model that is trained on data representing all the task difficulties, but if the decomposition of situational examples is done on the full data set $D_{i,t,c}$, then it may be possible to create dedicated models for each level of task difficulty (this choice depends on the actual possibility to sort task by difficulty at inference time, which may not be feasible for a variety of reasons).
Figure 34: Schematic representation of the process of learning and performance assessment of a skill program $Sol_{i,t,c,j}$, sophisticated by the distinct estimation of performance by level of task difficulty. In this more elaborated version, the set $E_{i,t,c[,q]}$ is decomposed in subsets of situational examples grouped by task difficulty $d$, such that the performance of the candidate model can be specifically measured by task difficulty.
This general treatment illustrates a procedural path from data to skill policy. However, while formally straightforward, an example inspired by a practical use case may contextualize even better the effort.
A.4.1 A practical example
To explore the logics and utility of task difficulty and human-machine augmentation as factors playing a role in the analysis of a skill policy decision problem, we walk through a practical example. It should be noted that the data presented in the following example is hypothetical and illustrative, not empirical, and has been constructed solely for clarifying purpose.
Context
The fictitious organization of interest for our example is a real-estate agency, whose business is based on commissions (3% of the transaction price). A core decision-making problem for the agency is to determine the optimal price for properties, such that the total commissions earned in a year is maximized. A price lower than the optimal one makes a sale faster, but reduces the satisfaction of the seller (the primary agency’s customer) and reduces the commission, which is computed as a percentage of the transaction price. In contrast, a price higher than the optimal one makes the sale slower and more difficult, increasing the total resources allocated by the agency to each specific sale opportunity and potentially reducing the number of transaction in the period of interest. As a result, the definition of an optimal price is a crucial factor for the agency to achieve economic health and wealth. Additionally, as the turnover of sales representatives is increasingly high, creating a proprietary machine skill that becomes a distinctive agency asset is seen as strategic from the agency’s shareholders. The mean cost of error is estimated as 10% of the MAPE (the true price for a property is computed by adjusting the actual transaction price by the lead time of sale, such that the true price tends to be higher than the true potential transaction price for quick sales, and tends to be lower in case of slow sales (the average transaction price in the last 3 years is 200k and the average number of sales per year is 150). In this context, the market of technological solutions is growing, and the agency has been offered two alternative solutions (A and B). Now, the decision-making problem faced by the agency is two-folded: first, deciding whether to decline both the models A and B, or introducing one of them; then, in case of the introduction of one or more model(s), defining the most economically-effective skill policy: automation of the prediction task (machine-exclusive skill policy), or human-machine interaction?
Task
Predict the value of a real-estate property ($y$) based on a set of attributes $X=\{x_{1},x_{2},\ldots,x_{p}\}$.
Skill model
$y=\hat{f}(X)$, whether deployed in the form of a heuristic (human skill) or a statistical model learned on data (machine skill).
Candidate human skills
Heuristics used by experienced sales agents, with knowledge of the market (both offer and demand). In this case, we consider two categories of agents: Senior (more than 5 years in the local business, or more than 100 sales), or Junior.
Candidate machine skills
Model A is an on-premise solution based on a linear regression, with a development fee of 200k and negligible operating costs. Model B is a cloud-based solution implementing an ensemble of regressors (linear, random forest, and neural network) with 50k development costs, licensed for 100k/year.
Scenarios
The agency plans to expanding its activity to new areas, with emerging new high-profile constructions but lower connections with the city center and fewer services. It is opinion of the agency director that predicting the prices in these new area may be more complex, because the target customer is different from the typical one know by the agency (then, higher levels of generalization difficulty are required in this context). Simplifying, we denote the prediction of prices for properties located in the typical area served by the agency as a low-difficulty task, while a prediction related to properties located in the new areas is considered a high-difficulty task. It is also opinion of the agency director that Senior sales representatives may be more effective in the former, while Junior reps may be more effective in the latter. To validate this conjecture, the agency director commissions an empirical analysis aimed at measuring the Mean Absolute Percentage Error (MAPE) of all the candidate predictive skills. The MAPE measured through the analysis is the following: 15% for Senior sales representative, 25% for the Junior ones, and 15% for both Model A and Model B.
Decision-making process
In a typical fashion, the predictive performance of the models and the corresponding economic impact are estimated for human-exclusive and machine-exclusive skill policy options without distinction by task difficulty (this is the measure captured by the metrics accompanying the models, as described in the Context). In Table 20 we illustrate the situation characterizing our example,with the models’ performance the related total cost. The total cost sums the cost of errors and the cost for development and operation. The total expected cost of error is computed multiplying the MAE, estimated through the MAPE, by the individual cost of error in percentage (10%, as specified in the Context). The incremental cost of skills is, for each period, the sum of the development cost and operational cost. For simplicity, the total cost is not actualized and the overall period of interest is 3 years. The optimization objective is to minimizing the total cost.
Table 20: Estimation of the economic impact (total cost) of human-exclusive and machine-exclusive skill policies over the entire task difficulty space. For example, in case of H_Senior: the cost of errors is equal to 200 * 150 * 0.10 * 0.15 = 450, where 200 is the average transaction price, 150 is the expected number of transactions per year, 0.10 is the fraction of the mean absolute error suffered as a cost, and 0.15 is the 15% MAPE for the skill H_Senior.
Under this informational scenario, the preferential choice for the economic-minded decision-maker is to allocate all the predictions to the human skills, specifically to Senior sales representatives. The machine skills, although achieving comparable execution performance and cost of errors, present higher total costs. The advantage of governing the execution of this task through a specific policy is remarkable, moving total cost from an average of 1,912.5 (representing the case of Senior and Junior sales reps being individually charged to make their own predictions) to 1,500 (representing the case of all predictions assigned to Senior reps), a whopping -21.6%.
Then, in Table 21 we see an example of the economic impact of the human and machines skills as seen through the specific lens of low-difficulty examples (properties in the typical area served by the agency) and high-difficulty examples (properties in the new areas). For machine skills, Model A and Model B are selectively trained and tested on low-difficulty and high-difficulty examples, creating at all effects four distinct models (as illustrated in the previous section, the test phase may require to consider multiple perturbations of the training/test set composition, but the final model is trained on all the examples available). For human skills it is not practically possible to compartmentalize the training phase on specific examples, so the models H_Senior and H_Junior are simply selectively tested on low/high-difficulty sets to determine more specific performance assessments with respect to Table 20.
Then, selectively deploying Model B (automation) on predictions expected to require low generalization difficulty, and the human-machine skill policy to high-difficulty predictions combining the skills of Junior sales reps with Model B, a total cost of 440 + 695 = 1,135 is achieved, reducing it of an additional 26.8% with respect to the H_Senior skill deployed as one-fits-all solution.
Crucially, while the combinations A+Senior provide low augmentation and B+Senior a moderate one, the combinations A+Junior and B+Junior present a visible and strong augmenting effect respectively, thus containing the cost of errors despite of the operational costs. As the data is not empirically gathered and constructed only for illustrative purposes, we cannot analyze from the bottom-up the specific reasons why these augmenting effects emerge. However, intuitively, this may happen when the models are extremely diverse in their ability to capture the nature of signals present in the data, thus making one model very effective in achieving high performance when the other systematically fails, and vice versa. In this case, the logic of interaction and cooperation between the models towards determining the final response is a critical factor for success, deserving a specific effort (for example, in a first instance, the relatively simple weighting of the responses provided by the models part of the $HM$ ensemble can be learned on data, after a reasonable period of dedicated data collection; in a further and more refined step of development, one may also think to find additional variables that allow to set specific weights for specific situations, thus creating a set of expert ensembles based upon standard elementary models).
Naturally, these results heavily depended on the specific point values chosen for the example, and do not represent a universal case. However, the order of magnitude of these values seems to match the common sense. As we can see, the values of MAPE with $d=Low$ and $d=High$ are consistent with the arithmetic average of the MAPE of the human-exclusive and machine-exclusive skill policies when measured without distinction of task difficulty (we assume uniform distribution of skills over tasks and uniform distribution of task difficulty).
Table 21: Estimation of the economic impact (total cost) of human-exclusive and machine-exclusive skill policies by task difficulty $d\in\{Low,High\}$.
In this example, simulation is by-passed by the possibility to estimate the economic impact of each skill policy option through a simple calculation based on MAPE, which is a synthetic central tendency measure, and not by a distribution of probabilities in the form of $P(APE)$. This simplified case is a specific instance of the more general process described in this section.